### v1

# Personalized Federated Learning with Heat-Kernel Enhanced Tensorized Multi-View Clustering

Abstract
We present a robust personalized federated learning framework that leverages heat-kernel enhanced tensorized multi-view fuzzy c-means clustering with advanced tensor decomposition techniques. Our approach integrates heat-kernel coefficients adapted from quantum field theory with Tucker decomposition and canonical polyadic decomposition (CANDECOMP/PARAFAC) to transform conventional distance metrics and efficiently represent high-dimensional multi-view structures. The framework employs matriculation and vectorization techniques to facilitate the discovery of hidden structures and multilinear relationships via N-way generalized tensors. The proposed method introduces a dual-level optimization scheme: local heat-kernel enhanced fuzzy clustering with tensor decomposition operating on order-N input tensors, and federated aggregation of tensor factors with privacy-preserving personalization mechanisms. The local stage employs tensorized kernel Euclidean distance transformations and Tucker decomposition to discover client-specific patterns in multi-view tensor data, while the global aggregation process coordinates tensor factors (core tensors and factor matrices) across clients through differential privacy-preserving protocols. This tensorized approach enables efficient handling of high-dimensional multi-view data with significant communication savings through low-rank tensor approximations.
Keywords: Federated learning, Multi-view clustering, Heat-kernel methods, Tensor decomposition, PARAFAC2, Tucker decomposition, Personalized machine learning, Privacy-preserving clustering.
1 Introduction
Large-scale data centers become increasingly important as the Internet of Things (IoT) continues to grow, with more devices being massively connected to the internet than ever before. This phenomenal transition of data to distributed environments affects the need for investigating optimal approaches to process heterogeneous information from multiple devices while ensuring security and efficient data exchange. Conventional unsupervised techniques cannot effectively handle these higher-order or higher-dimensional data structures inherent in modern federated systems where distributed clients possess different data modalities, feature dimensions, and local patterns.
The emergence of distributed multi-view data across heterogeneous clients presents unique challenges for federated learning systems. Traditional federated clustering assumes homogeneous data representations, failing to address the complexity of multi-view tensor structures where clients may have different modalities, feature spaces, and local data characteristics.
The massive growing of data is happening as the advance and revolution of technology in many fields becomes so forefront and powerful. Machine learning algorithms also developed and extending to significantly enables these data trends improved analysis. Mathematically, a data is representing by a unit or collection of row and column numbers arranged in a rectangular or array-like fashion. An array representing vectors or matrices storage set of different numbers we wish to have. In unsupervised clustering, we partition the rows unit data based on its given dimension attribute into some groups or clusters. This is a typical data that used as an input matrix to reveal important information in an efficient manner. However, a multi-resources data requires its representation to better process using tensor space aiming to meet the significant raise of data growth. Reconstruction tensor-based approaches to represent heterogeneous data can be challenging as it involves multilinear relationship between sets of matrix algebra but promising satisfaction output.
Most common application of third-order tensor representation is a color image. RGB image is stored in a 3-Dimensional array as $\text{channels}\times\text{height}\times\text{width}$. Fig. 1 displayed the illustration of R, G, and B component of images. As it can be seen, the shape of the image in Fig. 1 is $7\times 7\times 3$. The colour channels stored the height and width of the single image color of R, G, and B. Usually, the pixel values of a single image will be arranged in a 1-Dimensional vector to get the required format for classification or clustering process.
Figure 1: The Illustration of A 3-Dimensional of a Red-Green-Blue (RGB) Image
Tensor-based approaches had been extended in many machines learning techniques, including K-means clustering. Singular value decomposition (SVD) and principal component analysis (PCA) [1, 2] are two feature reduction techniques which frequently used to construct the data before starting classification or clustering process. Both SVD and PCA can process a data matrix which is well known as tensor of order two. On the other hand, multi-view (MV) data collects information from multiple sources. Each sources represent their feature component about one observation or object in a data. Mathematically, each source can be regarded as view. Accordingly, view represented by matrix space. For instance, each matrix on MV data may represent video recording, images, text, or audio recording’ feature component. A representation of these data sources possibly achieves good performances with tensor-based reconstruction as it can be reshaping based on their multicollinearity. In tensor space, MV data considered as a third-order tensor, with size $n\times d_{h}\times s$.
From the perspective of MV clustering, each component of feature-view stored as a two-order tensor. In this sense, if one MV data contains four views, then MV clustering algorithm will cluster the data by processing these four-two-Dimensional matrix based on their similarities. Most commonly, the similarities are measured by estimating the Euclidean distance between data points and cluster centers within/across views. The analogy of diversity and heterogeneity on multiple representations of the same information is precise [3]. For instance, online media such as CNN and FOX featuring a news about technology or culture based on different observation depending on their own perspective-based, context-based, purpose-based, etc. The different writing styles, image or video, and short description provided is unique (“heterogeneous”) and has semantic relationship. Hence, the representation to determine the space of variations to occur the unity might be useful if the structuring processes involves tensor-based reconstruction forming more smaller matrices and reduce the time complexity of conventional MV clustering algorithms.
The conventional k-means clustering not effective to directly explore structure of tensor space as it is required lots of memory. To fit into memory required processes, we will enforce the amount of data greatly using CANDECOMP (PARAFAC2). Our main motivation in using tensor-based Tucker decomposition approaches is to reveal hidden information that often happening in one data. The elaboration of these decomposition of PARAFAC2 and Tucker model will be also implemented into MV data directing its tensor transformation and modification of k-means objective function for processing higher-order tensor data.
2 Methodology
The integration of federated learning with multi-view clustering and tensor decomposition represents a convergence of several active research domains. This section reviews the foundational work and recent advances in federated learning, multi-view clustering, tensor decomposition methods, and their intersections, positioning our contributions within the broader landscape of distributed machine learning.
2.1 Federated Learning and Clustering
Federated learning, pioneered by McMahan et al. [4], enables collaborative machine learning across distributed clients without centralizing raw data. The seminal FedAvg algorithm established the foundation for aggregating local model updates through weighted averaging. However, traditional federated learning primarily focuses on supervised tasks, with federated clustering receiving comparatively less attention despite its practical importance in distributed data analytics.
Konečnỳ et al. [5] introduced federated optimization as a method to train a centralized model using decentralized data stored across numerous devices, emphasizing communication efficiency due to limited connectivity and privacy concerns. They proposed modifications to existing algorithms to adapt them to federated optimization settings. Li et al. [6] adopted the foundational ideas of federated optimization and extended them by integrating advanced techniques like gradient compression, acceleration, and variance reduction to address communication bottlenecks and improve the scalability of federated learning systems
Early work on federated clustering by Dennis et al. [7], Ghosh et al. [8], and Makhija et al. [9] demonstrated the feasibility of distributed k-means and hierarchical clustering. These approaches typically involve local clustering followed by global aggregation of cluster centers or assignments. However, these methods assume homogeneous data distributions and fail to address the inherent heterogeneity in federated environments.
Recent advances have introduced personalized federated clustering to address client heterogeneity. Caldarola et al. [10] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. [11] developed meta-learning approaches for personalized cluster initialization. These works, however, are limited to single-view data and do not leverage the rich multi-modal information available in modern federated systems.
Privacy-preserving aspects of federated clustering have been addressed through differential privacy mechanisms [12] and secure aggregation protocols [13]. Geyer et al. [14] demonstrated how differential privacy can be integrated into federated clustering while maintaining acceptable utility. However, these privacy mechanisms often come at the cost of clustering accuracy, particularly in high-dimensional settings.
The communication efficiency challenge in federated clustering has been tackled through gradient compression [15], sparsification techniques [16], and model quantization [17]. While these methods reduce communication overhead, they typically do not exploit the inherent structure of multi-view tensor data for more efficient representation and transmission.
2.2 Multi-View Clustering
Multi-view clustering leverages complementary information from multiple data representations to improve clustering performance. Early work by Bickel and Scheffer [18] established the theoretical foundations for multi-view learning, demonstrating superior performance when views provide complementary information about the underlying data structure.
Classical multi-view clustering approaches can be categorized into three main paradigms: co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as the work by Kumar et al. [19], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [20] and shared subspace clustering [21], project multiple views into a common representation space for unified clustering.
Graph-based multi-view clustering has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. [22] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. [23] developed parameter-free auto-weighted multiple graph learning. These methods construct view-specific graphs and learn optimal view weights for clustering.
Deep learning approaches have revolutionized multi-view clustering through end-to-end representation learning. Zhang et al. [24] introduced auto-encoder based multi-view clustering, while Zhu et al. [25] developed deep multi-view clustering networks with view correlation discovery. However, these methods typically require centralized data access and are not designed for federated environments.
Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. [26] and Wang et al. [27], introducing weighted aggregation schemes and view-specific fuzzification parameters. These soft clustering approaches provide uncertainty quantification but lack the scalability and privacy considerations required for federated deployment.
The mathematical foundations of multi-view clustering build upon single-view clustering formulations. In this study, we focus on soft clustering approaches that provide probabilistic cluster assignments, enabling better uncertainty quantification in federated environments. The objective function of standard fuzzy c-means (FCM) or soft clustering can be expressed as:
$J_{FCM}(U,A)=\sum_{i=1}^{n}\sum_{k=1}^{c}\mu_{ik}^{m}\|x_{i}-a_{k}\|^{2}$ (1)
where $U$ is the membership matrix of data sets $X$ in clusters $c$, AA is the cluster centers matrices of the $j$-th feature in clusters $c$, and $m$ is the fuzzifier to control the distribution of membership matrix $U$. Eq. 1 is designed to address data with a single representation, which is not compatible to solve a problem with multiple representations. To tackle this, $s$ number of views is introduced, such as indicator $h=1,\ldots,s$ enabled the representation of data X from $s$ views. In this way, the clustering problems based on multiple representation data $X^{h}$ can be solved by using the following equation:
$J_{MVC}(U^{h},A^{h})=\sum_{h=1}^{s}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{h})^{m}\|x_{i}^{h}-a_{k}^{h}\|^{2}$ (2)
As displayed in Eq. 2, the memberships of the $h$-th view data in $c$ clusters $\mu_{ik}^{h}$ is computed locally. Usually, to get a global memberships of data view $s$, we have an additional step called as aggregation of collected memberships across data views $s$. There are multiple ways to aggregate the memberships of $s$ data views, i.e., using the averagation approach or taking a product of memberships of the $i$-th data point across all views $s$ in clusters $c$ and divide it with the number of data points. If we consider a weight factor $v_{h}$, we can multiply each weight factors with the memberships and divide it by the number of data views, etc. As we focused on a client’ multi-view data, we simplify the computation for each view memberships by directly implementing the global memberships from an early iteration. In this way, we can modify Eq. 2 in the following way
$J_{{MVC}_{2}}(U^{*},A^{h})=\sum_{h=1}^{s}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}\|x_{i}^{h}-a_{k}^{h}\|^{2}$ (3)
Our goal is to provide a comprehensive mathematical objective function to robustly handle multi-view data distributed on multiple clients using tensor algebra, with non-linearity compatibilities. Thus, we replace the distance between data points and cluster centers across data views in Euclidean space with a kernel Euclidean distance (KED). We define the mathematical expression of KED in the following way:
$\text{KED}(x_{ij}^{h},a_{kj}^{h})=\bigg\{1-\text{exp}\bigg(-\sum_{j=1}^{d_{h}}\delta_{ij}^{h}(x_{ij}^{h}-a_{kj}^{h})^{2}\bigg)\bigg\}$ (4)
where $\delta_{ij}^{h}$ is a heat kernel coefficient (HKC) to control the Euclidean distance between data points $x_{ij}^{h}$ and cluster centers $a_{kj}^{h}$ across data views in kernel space. In this way, we can modify Eq. 3 in the following way
$J_{KMVC}(U^{*},A^{h})=\sum_{h=1}^{s}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}\text{KED}(x_{ij}^{h},a_{kj}^{h})$ (5)
subject to:
$\sum_{k=1}^{c}\mu_{ik}^{*}=1,\mu_{ik}^{*}\in[0,1]$
In Eq. 5, we treat all data views equally important, which is not efficient to solve multiple resources data. Since different views represents different feature behavior, it is necessary to assign a different weight $v_{h}$ to each data view $x_{i}^{h}$, with a normalization approach $\sum_{h=1}^{s}v_{h}=1$. By considering weight factors $v_{h}$, we can modify Eq. 5 in the following way.
$J_{E-KMVC}(V_{h},U^{*},A^{h})=\sum_{h=1}^{s}v_{h}^{\alpha}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}\bigg\{1-\text{exp}\big(-\sum_{j=1}^{d_{h}}\delta_{ij}^{h}(x_{ij}^{h}-a_{kj}^{h})^{2}\big)\bigg\}$ (6)
subject to:
$\sum_{k=1}^{c}\mu_{ik}^{*}=1,\mu_{ik}^{*}\in[0,1],\quad\sum_{h=1}^{s}v_{h}=1,v_{h}\in[0,1]$
where $\alpha$ is an exponent parameter to control the distribution of data views during the clustering processes.
2.3 Heat Kernel Methods and Quantum-Inspired Learning
Heat kernel methods have emerged as powerful tools for capturing local geometric structure in high-dimensional data spaces. Originally developed in differential geometry and quantum field theory, heat kernels provide a natural way to encode multi-scale geometric information through diffusion processes on manifolds.
The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. [28] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon [29] established theoretical foundations for diffusion-based data analysis.
In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona [30] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg [31] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. [32], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [33] and quantum neural networks [34].
The integration of heat kernel coefficients with clustering algorithms represents a natural extension of kernel methods to soft clustering scenarios. By incorporating heat diffusion properties, clustering algorithms can better capture local data manifold structure and improve robustness to noise and outliers.
2.4 Federated Multi-View Learning and Privacy Preservation
The intersection of federated learning and multi-view clustering presents unique challenges that are not adequately addressed by existing literature. Most federated learning approaches assume single-view data and homogeneous client populations, while multi-view clustering methods typically require centralized data access.
Recent work on federated multi-view learning includes approaches by Chen et al. [35] for federated multi-view representation learning and Liu et al. [36] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication.
Privacy preservation in federated multi-view settings requires careful consideration of information leakage across views and clients. Differential privacy mechanisms must be adapted to handle the multi-view tensor structure while maintaining clustering utility. Secure aggregation protocols need to be extended to handle tensor factor aggregation efficiently.
The communication efficiency challenge is particularly acute in federated multi-view scenarios due to the high-dimensional nature of multi-view tensor data. Existing compression techniques are not optimized for tensor structures and may destroy important multilinear relationships. This motivates the need for tensor-aware compression and aggregation schemes.
So, we will discuss the tensorization and activation of Eq. 6 in federated environments, addressing these open challenges through our proposed heat-kernel enhanced tensorized framework in the next sections.
2.5 Tensor Decomposition Methods
Tensor decomposition has emerged as a powerful paradigm for analyzing high-dimensional multi-modal data, providing compact representations that capture multilinear relationships inherent in complex datasets. The field has evolved from classical matrix factorization techniques to sophisticated tensor algebra frameworks capable of handling arbitrary N-way data structures.
Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang [37] and Harshman [38], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. The uniqueness property of CPD makes it particularly valuable for component analysis and factor identification in multi-way data.
Tucker decomposition, proposed by Tucker [39], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [40] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems.
Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [41] and Yuan and Zhang [42], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [43] provide resilience against outliers and corrupted entries.
Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [44] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [45], enabling decompositions with non-negativity constraints that often yield more interpretable factors.
The integration of tensor decomposition with machine learning has led to tensorized neural networks [46], tensor-based kernel methods [47], and multi-task learning frameworks [48]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power.
In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Zhou et al. [49] developed tensor-based multiway arrays for bid data sets, while Benjamin et al. [50] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.
A tensor is a higher-order array that represents multi-dimensional data structures. Scalar is a zero-order tensor; Vector is a 1-dimensional array, also known as tensor of order one and denoted by boldface lowercase letter, e.g., $\mathbf{x}$; A matrix $\mathbf{X}$ is a 2-way array or tensor of order two with two indices $(m,n)$, where $m$ indicates the row dimension and $n$ indicates the column dimension; While a tensor $\mathcal{X}$ is a generalization of matrices with three or more indices $(i,j,k,\ldots)$, considered as 3-way arrays or higher-way or N-way arrays (higher-order) tensors.
Notation and Mathematical Framework: Let $X=\{x_{1},\ldots,x_{n}\}$ be the collection of $n$ data points in $\mathbb{R}^{d}$ with $x_{i}=[x_{ij}]_{n\times d}$, $i=1,\ldots,n$, $j=1,\ldots,d$; $A=\{a_{1},\ldots,a_{c}\}$ be the set of cluster centers with $a_{k}=[a_{kj}]_{c\times d}$ and $k=1,\ldots,c$, and $U=\{\mu_{1},\ldots,\mu_{c}\}$ be the cluster membership matrix with $\mu_{k}=[\mu_{ik}]_{n\times c}$.
In the multi-view scenario, $h$ denotes the view index, $h=1,\ldots,s$, and the feature dimensionality of the $h$-th view is denoted as $d_{h}$. Thus, $X^{h}=\{x_{1}^{h},\ldots,x_{n}^{h}\}$ defines the data input $X$ in the $h$-th view with $x_{i}^{h}=[x_{ij}^{h}]_{n\times d_{h}}$. Hence, $A^{h}=\{a_{1}^{h},\ldots,a_{c}^{h}\}_{h=1}^{s}$ represents the cluster centers of the $h$-th view with $a_{k}^{h}=[a_{kj}^{h}]_{c\times d_{h}}$.
When $X\in\mathbb{R}^{n\times d_{h}\times s}$ represents multi-view data in tensor space, the notation $\mathcal{X}_{ijh}$ denotes the score or element of the $i$-th data point on the $j$-th feature component at the $h$-th view. Figure 2 illustrates the representation of multi-view data in tensor space. As clearly shown, multi-view data can be viewed as a 3-dimensional tensor $\mathcal{X}=\{x_{ijh}\}_{i=1,j=1,h=1}^{n,d_{h},s}$ or equivalently as a set of 2D matrices $\mathcal{X}=\{X^{1},\ldots,X^{s}\}$ with size $n\times d_{h}$.
Figure 2: The Illustration of MV Data in Tensor Space
Tensor Operations and Decomposition Preliminaries: The elements of $N$-way arrays can be reordered into matrices or vectors through matricization and vectorization operations. The matricization is reordering tensor’ elements into a matrix, while vectorization is reordering tensor’ elements into a vector. Canonical polyadic decomposition (CPD) or parallel factors (PARAFAC) [51, 37], Tucker decomposition (TD), and non-negative tensor factorization (NTF) are well-known approaches to reorder a tensor into a matrix or vector. PARAFAC or CPD holds multilinearity assumptions such as handle a sum of rank-one tensors by targeting the same lengths. In this sense, the original PARAFAC decomposition can reveal uncorrelated (orthogonal) and correlated (oblique) factors in system-variation data matrices, but failed to describe correlated factors when applied to the cross-product matrices computed from one data [52, 53]. PARAFAC2, on the other hand has the advantageous to handle irregular tensors by relaxing the canonical polyadic (CP) model [52]. PARAFAC2 enables the three-mode model, often implemented into subspace clustering algorithms taking advantageous of correlations between pair of factors. The PARAFAC2 normalize the generalization of PARAFAC to deal both with uncorrelated and correlated factors.
2.5.1 The Basic Background of Tensors
This subsection provides some definitions and concepts of tensor algebra, most of overviews and notations are borrowed from Kolda et al. [54].
Definition 1: The $N$-th order of tensor $\mathcal{A}$ is expressed as
$\mathcal{A}=[\mathcal{A}_{g_{1},g_{2},\ldots,g_{n}}]\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ (7)
where the $(g_{1},g_{2},\ldots,g_{n})$-th elements of the tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is containing $(\mathcal{A})_{g_{1},g_{2},\ldots,g_{n}}$ or $a_{g_{1},g_{2},\ldots,g_{n}}$ for indices $g=\{1,2,\ldots,G_{1}\}$, $g=\{1,2,\ldots,G_{2}\}$, and $g_{n}=\{1,2,\ldots,G_{N}\}$.
Definition 2: The outer product of $N$ vectors of rank-one $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is defined as
$\mathcal{A}=a^{(1)}\circ a^{(2)}\circ\ldots\circ a^{(N)}$ (8)
where $\circ$” represents the vector outer product. Fig. 3 presented the illustration of rank-one third-order tensor $\mathcal{A}=x\circ y\circ z$. The $(i,j,k)$ element of $\mathcal{A}$ is given by $a_{ijk}=x_{i}y_{j}z_{k}$. In terms of element wise, the 3-way tensor can be written as $a_{ijk}\approx\sum_{r=1}^{R}x_{ir}y_{jr}z_{kr}$.
Figure 3: The Illustration of Rank-One Third-Order Tensor $\mathcal{A}=x\circ y\circ z$
Definition 3: The $n$-mode matrix product of $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is denoted by $\mathcal{A}_{(n)}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n}\times G_{n+1}\times\ldots\times G_{N}}$ contains the element of $a_{g_{1},g_{2},\ldots,g_{n}}$ at $g_{n}$-th row and at $j$th column, denoted as
$j=1+\sum_{\ell=1,\ell\neq n}^{N}(g_{\ell}-1)\prod_{m=1,m\neq n}^{\ell-1}G_{m}$ (9)
Definition 4: The $n$-mode product between tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ and a matrix $L\in\mathbb{R}^{J_{n}\times G_{n}}$ is defined as $\mathcal{B}=\mathcal{A}\times_{n}L$, where the product of tensor $\mathcal{B}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n}\times G_{n+1}\times\ldots\times G_{N}}$. In terms of element wise, the $n$-mode product of tensor $\mathcal{B}$ can be computed as
$b_{g_{1}g_{2}\ldots g_{n-1}j_{n}g_{n+1}\ldots g_{N}}=\sum_{g_{1}=1}^{g_{n}}a_{g_{1}g_{2}\ldots g_{n}\ldots g_{N}}b_{j_{n}g_{n}}$ (10)
Definition 5: The tensor inner product between tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ and $\mathcal{B}\in\mathbb{R}^{J_{1}\times J_{2}\times\ldots\times J_{m}\times\ldots\times J_{M}}$ is denoted as
$\mathcal{C}=\textlangle\mathcal{A},\mathcal{B}\textrangle_{n,m}$ (11)
The element wise of tensor inner product $\mathcal{A}$ and $\mathcal{B}$ can be written explicitly as below
$c_{g_{1}g_{2}\ldots g_{n-1}g_{n+1}\ldots g_{N}j_{1}j_{2}\ldots j_{m-1}j_{m+1}\ldots j_{M}}=\sum_{q_{1}=1}^{q_{p}}a_{g_{1}g_{2}\ldots g_{n-1}g_{q}g_{n+1}\ldots g_{N}}b_{j_{1}j_{2}\ldots j_{m-1}j_{q}j_{m+1}\ldots j_{M}}$ (12)
where $i_{n}=j_{m}=q_{p}$ and the tensor $\mathcal{C}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n+1}\times\ldots\times G_{N}\times J_{1}\times J_{2}\times\ldots\times J_{m-1}\times J_{m+1}\times\ldots\times J_{M}}$.
Definition 6: The tensor scalar product of tensors $\mathcal{A},\mathcal{B}\in\coprod^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n+1}\times\ldots\times G_{N}}$ is the inner product of tensor $\mathcal{A}$ and $\mathcal{B}$, denoted as below
$\langle{\mathcal{A},\mathcal{B}}\rangle\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\sum_{g_{1}=1}^{G_{1}}\sum_{g_{2}=1}^{G_{2}}\ldots\sum_{g_{n}=1}^{G_{N}}a_{g_{1}g_{2}\ldots g_{n-1}g_{n+1}\ldots g_{N}}b_{g_{1}g_{2}\ldots g_{n-1}g_{n+1}\ldots g_{N}}$ (13)
Definition 7: The Frobenius norm of tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ also known as the “energy” of the tensor, defined as
$\|\mathcal{A}\|_{F}=\varepsilon_{\mathcal{A}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\sqrt{\langle{\mathcal{A},\mathcal{A}}\rangle}$ (14)
2.5.2 The Higher Order Singular Value Decomposition (HOSVD)
Tensor factorization had been implemented as a decomposition-based approach to project a large amount of data into a small number of data considering collinearity of its heterogeneous [55]. CANDECOMP/PARAFAC is one of the most existing tensor decomposition techniques to improve data clustering processes [54, 56]. The HOSVD is a multidimensional extension to the matrix singular value decomposition (SVD). To better explain the higher-order SVD, we will borrow some concept of Wang and Song [55] with a two-mode of tensor order $N$.
2.5.3 Tensor Distance
We know that, the Euclidean distance between data points $x_{i}$ and cluster centers $a_{k}$ can be represented as:
$d(x_{i},a_{k})=\|x_{i}-a_{k}\|^{2}$ (15)
Definition 8: The tensor distance of a given $\mathcal{X},\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{N}}$ $(N\geq 1)$, and $x$, $a$ denoted as the vector representation of tensor $\mathcal{X},\mathcal{A}$, such as the element wise of tensor $\mathcal{XA}_{G_{1}\times G_{2}\times\ldots\times G_{N}}$ $(1\leq g_{i}\leq G_{n},1\leq i\leq N)$ in $\mathcal{X},\mathcal{A}$ is corresponding to $x_{l}$, $a_{l}$, i.e., the $l$-th element in $x$, the $l$-th element in $a$, where $l=1+\sum_{i^{\prime}=1,i^{\prime}\neq i}^{G_{1}\times G_{2}\times\ldots\times G_{N}}(g_{l}-1)\prod_{m=1,m\neq i}^{i^{\prime}-1}G_{m}$. Then the tensor distance (TD) between $\mathcal{X}$ and $\mathcal{A}$ can be defined as
$d_{TD}(\mathcal{X},\mathcal{A})=\sqrt{\sum_{l,m=1}^{G_{1}\times G_{2}\times\ldots\times G_{N}}\rho_{lm}(x_{l}-a_{l})(x_{m}-a_{m})}=\sqrt{(x-a)^{T}P(x-a)}$ (16)
where $\rho_{lm}$ is the metric coefficient defined by
$\rho_{lm}=\frac{1}{2\pi\theta}\frac{\text{ exp }\big\{\|q_{1}-q_{m}\|_{2}^{2}\big\}}{2\theta}$ (17)
and $\|q_{1}-q_{m}\|_{2}$ is defined as the location distance between $\mathcal{X}_{g_{1}g_{2}\ldots g_{N}}$ and $\mathcal{A}_{g_{1}g_{2}\ldots g_{N}}$
$\|q_{1}-q_{m}\|_{2}=\sqrt{(g_{1}+g_{1}^{\prime})^{2}+(g_{2}+g_{2}^{\prime})^{2}+\ldots+(g_{n}+g_{n}^{\prime})^{2}+\ldots+(g_{N}+g_{N}^{\prime})^{2}}$ (18)
Here, $\theta$ is a user-defined parameter with $\theta>0$ and $P$ is the metric matrix.
2.6 Summary and Research Gaps
The review of existing literature reveals several important research gaps that motivate our proposed approach:
Limited Integration: While federated learning, multi-view clustering, and tensor decomposition have been extensively studied independently, their integration remains largely unexplored. Existing federated clustering methods do not leverage multi-view structure, and multi-view clustering approaches are not designed for federated environments.
Communication Inefficiency: Current federated multi-view methods do not exploit tensor structure for communication compression. The high-dimensional nature of multi-view data exacerbates communication overhead in federated settings, requiring novel tensor-aware compression schemes.
Privacy Limitations: Existing privacy-preserving mechanisms are not optimized for multi-view tensor data, potentially leading to information leakage across views or reduced clustering utility. New differential privacy mechanisms tailored to tensor structures are needed.
Lack of Personalization: Most federated clustering approaches assume homogeneous data distributions across clients. In reality, clients often have heterogeneous data characteristics requiring personalized clustering solutions that adapt to local data patterns while benefiting from global knowledge.
Theoretical Understanding: The convergence properties and privacy guarantees of federated multi-view tensor clustering are not well understood. Theoretical analysis is needed to provide convergence guarantees and characterize the privacy-utility trade-offs.
Our work addresses these gaps by proposing a novel heat-kernel enhanced tensorized framework that integrates advanced tensor decomposition with personalized federated learning, providing both theoretical guarantees and practical advantages in communication efficiency and clustering performance.
3 Problem Formulation
Let $\mathcal{L}=\{L_{1},L_{2},\ldots,L_{M}\}$ be a set of $M$ clients in the federated system. Each client $L_{\ell}$ possesses local multi-view data that can be organized into tensor form $\mathcal{X}_{[\ell]}\in\mathbb{R}^{n(\ell)\times d_{1}^{(\ell)}\times d_{2}^{(\ell)}\times\cdots\times d_{s(\ell)}^{(\ell)}}$, where:
  • $n(\ell)$ is the number of samples at client $\ell$
  • $s(\ell)$ is the number of views at client $\ell$
  • $d_{h}^{(\ell)}$ is the feature dimension of view $h$ at client $\ell$
Alternatively, we can represent the data as a collection of view-specific matrices $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ where $X_{[\ell]}^{h}\in\mathbb{R}^{n(\ell)\times d_{h}^{(\ell)}}$, which can be tensorized for advanced decomposition techniques.
3.1 Federated Kernel Euclidean Distance
In federated environment, we assume participated clients shared different number of data but the same feature component. Thus, the ground truth or predicted clusters of multi-view data sets stored on their data centers may vary. In this way, each data points on clients will shape a unified global multi-view datasets in a certain problem or case. Therefore, it is explained a lot the definition of “heterogeneous” on data. This characteristic is denoted by $c(\ell)$, representing the number of clusters of multi-view data sets held by the $\ell$-th client. By following this direction, we extend the heat-kernel enhanced distance metric KED in Eq. 4 to operate in federated environment. For client $\ell$, the kernel Euclidean distance (KED) between $x_{[\ell]ij}^{h}$ (representing data point $i$, $\forall i=1,\ldots,n(\ell)$) and cluster center tensor $a_{[\ell]kj}^{h}$ (representing cluster $k$, $\forall k=1,\ldots,c(\ell)$) is:
$\text{FKED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})=1-\exp\left(-\sum_{h=1}^{s(\ell)}\sum_{j=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij}^{h}\left(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h}\right)^{2}\right)$ (19)
where $\delta_{[\ell]ij}^{h}$ are heat-kernel coefficients held by $M$ clients.
3.2 Federated Heat-Kernel Coefficients
We focus on developing federated heat-kernel coefficients (FedH-KC), which are computed locally by each client to transform feature distances into privacy-preserving exponential kernel measures tailored to the federated setting. For the $j$-th feature in view $h$, we define FedH-KC $\delta_{\left[\ell\right]ij}^{h}$ using two distinct estimators.
$\text{FedH-KC}=\delta_{\left[\ell\right]ij}^{h}=\frac{x_{\left[\ell\right]ij}^{h}-\min_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)}{\max_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)-\min_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)+\epsilon}\qquad\forall j,h,\ell$ (20)
where $\epsilon>0$ prevents division by zero when features have identical values. This normalizes features within each view for consistent distance calculations across clients with different data distributions. and
$\text{FedH-KC}=\delta_{\left[\ell\right]ij}^{h}=\left|{x_{\left[\ell\right]ij}^{h}-{{\bar{x}}_{\left[\ell\right]j}}^{h}}\right|\qquad\forall j,h,\ell$ (21)
where $\bar{x}_{[\ell]j}^{h}=\frac{1}{n(\ell)}\sum_{i=1}^{n(\ell)}x_{[\ell]ij}^{h}$ represents the mean value of the $j$-th feature in view $h$ for client $\ell$, harmonizing with the centralized case notation.
3.3 Personalized Federated Learning
Considers there are $M$ number of clients storing multi-view data sets $X_{[\ell]}=X_{[\ell]i}^{h}$ for all $\ell=1,\ldots,M$ and $h=1,\ldots,s(\ell)$. Here, we assume each client is storing a different number of dimensionality of data points $x_{i}$. To solve clustering problem for these multi-view data sets distributed on $M$ clients, we propose an efficient federated kernel multi-view clustering (E-FKMVC). The proposed E-FKMVC modified Eq. 6 in the following way
$J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})=\sum_{\ell=1}^{M}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{~FKED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})$ (22)
subject to:
$\sum_{k=1}^{c}\mu_{[\ell]ik}^{*}=1,\mu_{[\ell]ik}^{*}\in[0,1]$ (23)
$\sum_{h=1}^{s}v_{[\ell]h}=1,v_{[\ell]h}\in[0,1]$ (24)
Due to the non-convex nature of the E-FKMVC objective function in Eq. 22, direct minimization is not feasible. We employ an alternating optimization strategy that iteratively updates the membership matrix $U_{[\ell]}^{*}$, view weights $V_{[\ell]}$, and cluster centers $A_{[\ell]}^{h}$ until convergence.
To derive the update rules, we formulate the Lagrangian incorporating the normalization constraints:
$\begin{split}\tilde{J}_{E-FKMVC}^{\ell}&=J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})\\ &\quad+\sum_{i=1}^{n(\ell)}{\lambda_{1i}}\left(\sum_{k=1}^{c(\ell)}{\mu_{[\ell]ik}^{*}-1}\right)+{\lambda_{2}}\left(\sum_{h=1}^{s(\ell)}{v_{[\ell]h}-1}\right)\end{split}$ (25)
where $\lambda_{1i}$} and $\lambda_{2}$} are Lagrange multipliers enforcing the fuzzy clustering constraints: membership normalization and view weight normalization, respectively.
Theorem 1 (E-FKMVC Update Rules).
For the federated objective function $J_{E-FKMVC}$ defined in Eq. 22, the necessary conditions for optimality yield the following update rules for client $\ell$:
Membership Matrix Update:
$\mu_{\left[\ell\right]ik}^{*}=\frac{{{{\left({\sum\limits_{h=1}^{s\left(\ell\right)}{v_{\left[\ell\right]h}^{\alpha}{\rm{FKED}}\left({x_{\left[\ell\right]ij}^{h},a_{\left[\ell\right]kj}^{h}}\right)}}\right)}^{-{{\left({m-1}\right)}^{-1}}}}}}{{\sum\limits_{k^{\prime}=1}^{c\left(\ell\right)}{{{\left({\sum\limits_{h=1}^{s\left(\ell\right)}{v_{\left[\ell\right]h}^{\alpha}{\rm{FKED}}\left({x_{\left[\ell\right]ij}^{h},a_{\left[\ell\right]k^{\prime}j}^{h}}\right)}}\right)}^{-{{\left({m-1}\right)}^{-1}}}}}}}$ (26)
Cluster Centers Update:
$a_{\left[\ell\right]kj}^{h}=\frac{{\sum\limits_{i=1}^{n\left(\ell\right)}{{{\left({\mu_{\left[\ell\right]ik}^{*}}\right)}^{m}}v_{\left[\ell\right]h}^{\alpha}\exp\left({-\delta_{\left[\ell\right]ij}^{h}{{\left\|{x_{\left[\ell\right]i}^{h}-a_{\left[\ell\right]k}^{h}}\right\|}^{2}}}\right)}}}{{\sum\limits_{i=1}^{n\left(\ell\right)}{{{\left({\mu_{\left[\ell\right]ik}^{*}}\right)}^{m}}v_{\left[\ell\right]h}^{\alpha}\exp\left({-\delta_{\left[\ell\right]ij}^{h}{{\left\|{x_{\left[\ell\right]i}^{h}-a_{\left[\ell\right]k}^{h}}\right\|}^{2}}}\right)}}}x_{\left[\ell\right]ij}^{h}$ (27)
View Weights Update:
${v_{\left[\ell\right]h}}=\frac{{{{\left({\sum\limits_{i=1}^{n\left(\ell\right)}{\sum\limits_{k=1}^{c\left(\ell\right)}{{{\left({\mu_{\left[\ell\right]ik}^{*}}\right)}^{m}}{\rm{FKED}}\left({x_{\left[\ell\right]ij}^{h},a_{\left[\ell\right]kj}^{h}}\right)}}}\right)}^{-{{\left({\alpha-1}\right)}^{-1}}}}}}{{\sum\limits_{h^{\prime}=1}^{s\left(\ell\right)}{{{\left({\sum\limits_{i=1}^{n\left(\ell\right)}{\sum\limits_{k=1}^{c\left(\ell\right)}{{{\left({\mu_{\left[\ell\right]ik}^{*}}\right)}^{m}}{\rm{FKED}}\left({x_{\left[\ell\right]ij}^{h^{\prime}},a_{\left[\ell\right]kj}^{h^{\prime}}}\right)}}}\right)}^{-{{\left({\alpha-1}\right)}^{-1}}}}}}}$ (28)
where the updates are performed iteratively until convergence, subject to the normalization constraints in Eqs. 23 and 24.
Proof.
We establish the necessary optimality conditions for the E-FKMVC objective function using Lagrangian optimization and derive the closed-form update rules for each parameter set. The proof proceeds by applying the method of Lagrange multipliers to the constrained optimization problem and solving the resulting system of equations.
Part I: Membership Matrix Update Rule Derivation
To derive the update equation for membership coefficients $\mu_{[\ell]ik}^{*}$ for client $\ell$, we begin by computing the partial derivative of the Lagrangian in Eq. 25 with respect to $\mu_{[\ell]ik}^{*}$.
Taking the partial derivative with respect to $\mu_{[\ell]ik}^{*}$:
$\frac{\partial\tilde{J}_{E-FKMVC}^{\ell}}{\partial\mu_{[\ell]ik}^{*}}$ $=\frac{\partial}{\partial\mu_{[\ell]ik}^{*}}\left[\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right]+\lambda_{1i}$ (29)
$=m(\mu_{[\ell]ik}^{*})^{m-1}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})+\lambda_{1i}$ (30)
Setting this derivative equal to zero for optimality:
$m(\mu_{[\ell]ik}^{*})^{m-1}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})+\lambda_{1i}=0$ (31)
Solving for $\mu_{[\ell]ik}^{*}$:
$(\mu_{[\ell]ik}^{*})^{m-1}$ $=-\frac{\lambda_{1i}}{m\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})}$ (32)
$\mu_{[\ell]ik}^{*}$ $=\left(-\frac{\lambda_{1i}}{m}\right)^{1/(m-1)}\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(m-1)}$ (33)
To determine the Lagrange multiplier $\lambda_{1i}$, we apply the normalization constraint $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$:
$\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}$ $=\left(-\frac{\lambda_{1i}}{m}\right)^{1/(m-1)}\sum_{k=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(m-1)}=1$ (34)
Solving for the Lagrange multiplier:
$\left(-\frac{\lambda_{1i}}{m}\right)^{1/(m-1)}$ $=\frac{1}{\sum_{k^{\prime}=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k^{\prime}}^{h})\right)^{-1/(m-1)}}$ (35)
Substituting this back into Eq. 33:
$\mu_{[\ell]ik}^{*}=\frac{\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(m-1)}}{\sum_{k^{\prime}=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k^{\prime}}^{h})\right)^{-1/(m-1)}}$ (36)
This establishes the membership matrix update rule in Eq. 26.
Part II: Cluster Centers Update Rule Derivation
For cluster center updates $a_{[\ell]kj}^{h}$, we differentiate the objective function in Eq. 22 with respect to $a_{[\ell]kj}^{h}$:
$\frac{\partial J_{E-FKMVC}^{\ell}}{\partial a_{[\ell]kj}^{h}}$ $=\frac{\partial}{\partial a_{[\ell]kj}^{h}}\left[v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right]$ (37)
$=v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\frac{\partial\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})}{\partial a_{[\ell]kj}^{h}}$ (38)
To compute $\frac{\partial\text{FKED}}{\partial a_{[\ell]kj}^{h}}$, we recall that:
$\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})=1-\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$ (39)
Let $\phi_{[\ell]ik}^{h}=\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}$. Then:
$\frac{\partial\text{FKED}}{\partial a_{[\ell]kj}^{h}}$ $=\frac{\partial}{\partial a_{[\ell]kj}^{h}}\left[1-\exp(-\phi_{[\ell]ik}^{h})\right]$ (40)
$=\exp(-\phi_{[\ell]ik}^{h})\frac{\partial\phi_{[\ell]ik}^{h}}{\partial a_{[\ell]kj}^{h}}$ (41)
$=\exp(-\phi_{[\ell]ik}^{h})\cdot 2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\cdot(-1)$ (42)
$=-2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$ (43)
Substituting this into Eq. 38 and setting equal to zero for optimality:
$v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\left[-2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)\right]=0$ (44)
Simplifying and rearranging:
$\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\delta_{[\ell]ij}^{h}$ $\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})=0$ (45)
Let $w_{[\ell]ijk}^{h}=(\mu_{[\ell]ik}^{*})^{m}v_{[\ell]h}^{\alpha}\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$. Then:
$\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})$ $=0$ (46)
$\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}x_{[\ell]ij}^{h}$ $=a_{[\ell]kj}^{h}\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}$ (47)
$a_{[\ell]kj}^{h}$ $=\frac{\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}x_{[\ell]ij}^{h}}{\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}}$ (48)
This yields the cluster center update rule in Eq. 27.
Part III: View Weights Update Rule Derivation
For view weight updates $v_{[\ell]h}$, we differentiate the Lagrangian with respect to $v_{[\ell]h}$:
$\frac{\partial\tilde{J}_{E-FKMVC}^{\ell}}{\partial v_{[\ell]h}}$ $=\frac{\partial}{\partial v_{[\ell]h}}\left[\sum_{h^{\prime}=1}^{s(\ell)}v_{[\ell]h^{\prime}}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h^{\prime}},a_{[\ell]k}^{h^{\prime}})\right]+\lambda_{2}$ (49)
$=\alpha v_{[\ell]h}^{\alpha-1}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})+\lambda_{2}$ (50)
Setting this derivative equal to zero for optimality:
$\alpha v_{[\ell]h}^{\alpha-1}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})+\lambda_{2}=0$ (51)
Solving for $v_{[\ell]h}$:
$v_{[\ell]h}^{\alpha-1}$ $=-\frac{\lambda_{2}}{\alpha\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})}$ (52)
$v_{[\ell]h}$ $=\left(-\frac{\lambda_{2}}{\alpha}\right)^{1/(\alpha-1)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(\alpha-1)}$ (53)
Applying the normalization constraint $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$:
$\sum_{h=1}^{s(\ell)}v_{[\ell]h}$ $=\left(-\frac{\lambda_{2}}{\alpha}\right)^{1/(\alpha-1)}\sum_{h=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(\alpha-1)}=1$ (54)
Solving for the Lagrange multiplier:
$\left(-\frac{\lambda_{2}}{\alpha}\right)^{1/(\alpha-1)}$ $=\frac{1}{\sum_{h^{\prime}=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h^{\prime}},a_{[\ell]k}^{h^{\prime}})\right)^{-1/(\alpha-1)}}$ (55)
Substituting back:
$v_{[\ell]h}=\frac{\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(\alpha-1)}}{\sum_{h^{\prime}=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h^{\prime}},a_{[\ell]k}^{h^{\prime}})\right)^{-1/(\alpha-1)}}$ (56)
This establishes the view weight update rule in Eq. 28.
Part IV: Convergence Analysis and Optimality Conditions
The derived update rules satisfy the Karush-Kuhn-Tucker (KKT) conditions for the constrained optimization problem. Specifically:
  1. 1.
    Stationarity: $\nabla J_{E-FKMVC}^{\ell}+\sum_{i=1}^{n(\ell)}\lambda_{1i}\nabla g_{i}+\lambda_{2}\nabla h=0$, where $g_{i}(\mu)=\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}-1$ and $h(v)=\sum_{h=1}^{s(\ell)}v_{[\ell]h}-1$.
  2. 2.
    Primal feasibility: $g_{i}(\mu)=0$ and $h(v)=0$ for all $i$.
  3. 3.
    Dual feasibility: All Lagrange multipliers are finite and well-defined.
The alternating optimization scheme converges to a local optimum under the following conditions:
  • Boundedness and Continuity: The objective function $J_{E-FKMVC}^{\ell}$ is continuous and bounded below. This follows from the fact that FKED values are bounded in $[0,1]$ by construction (Eq. 4), ensuring the weighted sum in the objective function remains finite. The exponential decay in the heat kernel guarantees that distances approach finite limits as data points move infinitely far apart.
  • Constraint Compactness: The constraint set is compact and convex. The membership matrix constraints $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$ with $\mu_{[\ell]ik}^{*}\in[0,1]$ define probability simplices, which are compact convex sets in $\mathbb{R}^{c(\ell)}$. Similarly, the view weight constraints $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$ with $v_{[\ell]h}\in[0,1]$ form a compact convex simplex in $\mathbb{R}^{s(\ell)}$. The cluster center parameters are unconstrained but remain bounded due to the data distribution.
  • Unique Subproblem Solutions: Each subproblem (updating $U$, AA, or $V$ while fixing others) has a unique solution. For membership updates, the strictly positive denominators in Eq. 26 ensure uniqueness since FKED values are always positive for distinct data points and cluster centers. For cluster center updates, the weighted exponential terms in Eq. 27 create a strictly convex weighted least-squares problem with a unique minimum. For view weight updates, the strictly decreasing nature of the power function with exponent $1/(\alpha-1)$ in Eq. 28 guarantees uniqueness of the normalized solution.
These conditions are satisfied by construction, as the FKED function is continuous and bounded, the probability simplex constraints define compact convex sets, and the update rules yield unique solutions when the denominators are non-zero (guaranteed by the strict positivity of FKED values and the non-degeneracy assumption that no data points are identical to cluster centers).
3.4 Tensorized Federated Objective
Based on Eq. 22, we propose the global objective combines local tensorized mullti-view clustering with federated coordination (T-MVCF):
$\min_{\{\mathcal{U}_{[\ell]},\mathcal{A}_{[\ell]},\mathcal{V}_{[\ell]}\},\{\mathcal{A}_{g},\mathcal{V}_{g}\}}\sum_{\ell=1}^{M}\left[\omega_{[\ell]}\mathcal{L}_{tensor}^{(\ell)}(\mathcal{X}_{[\ell]},\mathcal{U}_{[\ell]},\mathcal{A}_{[\ell]},\mathcal{V}_{[\ell]})+\gamma\mathcal{R}_{tensor}(\mathcal{A}_{[\ell]},\mathcal{A}_{g})+\eta\mathcal{R}_{view}(\mathcal{V}_{[\ell]},\mathcal{V}_{g})\right]$ (57)
where:
  • $\mathcal{L}_{tensor}^{(\ell)}$ is the local tensorized heat-kernel enhanced clustering loss
  • $\mathcal{U}_{[\ell]}$ represents the membership tensor for client $\ell$
  • $\mathcal{A}_{[\ell]}$ represents the tensor of cluster centers for client $\ell$
  • $\mathcal{V}_{[\ell]}$ represents the view weight tensor for client $\ell$
  • $\mathcal{R}_{tensor}$ and $\mathcal{R}_{view}$ are tensorized regularization terms ensuring global coordination
3.5 Tensor Decomposition Framework
To handle the high-dimensional tensor structures efficiently, we employ tensor decomposition techniques. For client $\ell$, we decompose the cluster center tensor using Tucker decomposition:
$\mathcal{A}_{[\ell]}\approx\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}^{(1)}\times_{2}\mathbf{U}_{[\ell]}^{(2)}\times_{3}\mathbf{U}_{[\ell]}^{(3)}$ (58)
where $\mathcal{G}_{[\ell]}$ is the core tensor and $\mathbf{U}_{[\ell]}^{(z)}$ are the factor matrices for mode $z$. This decomposition enables:
  1. 1.
    Dimensionality reduction through low-rank approximation
  2. 2.
    Efficient storage and communication of tensor parameters
  3. 3.
    Capture of multilinear relationships across different data modes
4 Proposed Federated Tensorized Clustering
Our proposed personalized federated learning framework integrates heat-kernel enhanced multi-view clustering with tensorized decomposition techniques at the local client level and coordinated global aggregation. The framework consists of three main components: (1) local tensorized heat-kernel enhanced multi-view fuzzy clustering, (2) tensor decomposition for efficient representation, and (3) federated aggregation with personalization mechanisms.
4.1 Tensorized Heat-Kernel Enhanced Local Clustering
Each client $L_{\ell}$ possesses multi-view data $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ where $X_{[\ell]}^{h}\in\mathbb{R}^{n(\ell)\times d_{h}^{(\ell)}}$ represents the $h$-th view. We tensorize this data structure to capture multilinear relationships across samples, features, and views.
4.1.1 Tensor Data Organization
The multi-view data at client $\ell$ is organized into a 3-mode tensor:
$\mathcal{X}_{[\ell]}\in\mathbb{R}^{n(\ell)\times D_{[\ell]}\times s(\ell)}$ (59)
where $D_{[\ell]}=\max_{h}d_{h}^{(\ell)}$ is the maximum feature dimension (with zero-padding for views with fewer features). This tensorization enables:
  • Unified representation across heterogeneous views
  • Capture of cross-view feature interactions
  • Efficient tensor decomposition operations
4.1.2 Tensorized Kernel Euclidean Distance
We extend the heat-kernel enhanced distance metric to operate on tensor slices. For client $\ell$, the tensorized kernel Euclidean distance between data tensor slice $\mathcal{X}_{[\ell],i,:}^{h}$ (representing data point $i$) and cluster center tensor $\mathcal{A}_{[\ell],k,:}^{h}$ (representing cluster $k$) is:
$\text{TKED}_{\ell}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})=1-\exp\left(-\sum_{h=1}^{s(\ell)}\sum_{j=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij}^{h}\left(\mathcal{X}_{[\ell],i,j}^{h}-\mathcal{A}_{[\ell],k,j}^{h}\right)^{2}\right)$ (60)
where $\delta_{[\ell]ij}^{h}$ are tensorized heat-kernel coefficients held by the $\ell$-th client, computed as:
$\delta_{[\ell]ij}^{h}=\frac{\mathcal{X}_{[\ell],i,j}^{h}-\min_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}}{\max_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}-\min_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}}$ (61)
4.1.3 Tensor Decomposition for Cluster Centers
To reduce computational complexity and enable efficient communication, we decompose the cluster center tensor using Tucker decomposition:
$\mathcal{A}_{[\ell]}\approx\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$ (62)
where:
  • $\mathcal{G}_{[\ell]}\in\mathbb{R}^{c\times r_{2}\times r_{3}}$ is the core tensor
  • $\mathbf{P}_{[\ell]}\in\mathbb{R}^{c\times c}$ captures cluster relationships
  • $\mathbf{Q}_{[\ell]}\in\mathbb{R}^{D_{[\ell]}\times r_{2}}$ captures feature patterns
  • $\mathbf{R}_{[\ell]}\in\mathbb{R}^{s(\ell)\times r_{3}}$ captures view relationships
The decomposed representation reduces storage from $\mathcal{O}(c\cdot D_{[\ell]}\cdot s(\ell))$ to $\mathcal{O}(c\cdot r_{2}\cdot r_{3}+c^{2}+D_{[\ell]}\cdot r_{2}+s(\ell)\cdot r_{3})$.
4.1.4 Tensorized Local Objective Function
Each client optimizes a tensorized multi-view fuzzy clustering (T-MVFC) objective:
$J_{\ell}^{tensor}(\mathcal{U}_{[\ell]},\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]},\mathbf{v}_{[\ell]})=\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h}^{\alpha}\text{ TKED}_{\ell}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})$ (63)
subject to the constraints:
$\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}$ $=1,\quad\mu_{[\ell]ik}^{*}\in[0,1]$ (64)
$\sum_{h=1}^{s(\ell)}v_{[\ell]h}$ $=1,\quad v_{[\ell]h}\in[0,1]$ (65)
$\mathcal{A}_{[\ell]}$ $=\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$ (66)
Theorem 2 (Tensorized Update Rules).
The alternating optimization of the tensorized objective function yields:
Membership updates:
$\mu_{jk}^{(i)}=\frac{\left(\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\text{TKED}_{i}(j,k)\right)^{-\frac{1}{m-1}}}{\sum_{k^{\prime}=1}^{c}\left(\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\text{TKED}_{i}(j,k^{\prime})\right)^{-\frac{1}{m-1}}}$ (67)
Core tensor updates:
$\mathcal{G}_{i}=\arg\min_{\mathcal{G}}\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\|\mathcal{X}_{i,j,:,h}-(\mathcal{G}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i})_{k,:,h}\|^{2}$ (68)
Factor matrix updates:
$\mathbf{P}_{i}$ $=\arg\min_{\mathbf{P}}\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\|\mathcal{X}_{i,j,:,h}-(\mathcal{G}_{i}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i})_{k,:,h}\|^{2}$ (69)
$\mathbf{Q}_{i}$ $=\arg\min_{\mathbf{Q}}\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\|\mathcal{X}_{i,j,:,h}-(\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}\times_{3}\mathbf{R}_{i})_{k,:,h}\|^{2}$ (70)
$\mathbf{R}_{i}$ $=\arg\min_{\mathbf{R}}\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\|\mathcal{X}_{i,j,:,h}-(\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R})_{k,:,h}\|^{2}$ (71)
View weight updates:
$v_{h}^{(i)}=\frac{\left(\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\text{TKED}_{i}(j,k,h)\right)^{-\frac{1}{\alpha-1}}}{\sum_{h^{\prime}=1}^{s_{i}}\left(\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\text{TKED}_{i}(j,k,h^{\prime})\right)^{-\frac{1}{\alpha-1}}}$ (72)
Proof.
The proof follows similar principles to the non-tensorized case but operates on tensor slices. The factor matrix updates can be solved using alternating least squares (ALS) with weighted terms based on membership and view weights. ∎
4.1.5 Parallel Local Clustering Algorithm
Algorithm 1 Efficient Federated Kernel Multi-View Clustering (E-FKMVC)
Input: Local multi-view data $X_{i}$, global centers $A_{g}$, personalization parameter $\lambda_{i}$
Initialize: $A_{i}^{(0)}\leftarrow\lambda_{i}A_{g}+(1-\lambda_{i})A_{i}^{local}$, $V_{i}^{(0)}\leftarrow\frac{1}{s_{i}}\mathbf{1}$
$t\leftarrow 0$
repeat
  $t\leftarrow t+1$
  // Compute heat-kernel coefficients
  Calculate $\delta_{j\ell}^{h,(i)}$ using Eq. 20 or 21
  // Update memberships
  Update $U_{i}^{(t)}$ using Eq. 26
  // Update cluster centers
  Update $A_{i}^{(t)}$ using Eq. 27
  // Update view weights
  Update $V_{i}^{(t)}$ using Eq. 28
  // Check convergence
  Compute $J_{i}^{local}$ using Eq. 22
until convergence or max iterations
return $U_{i}^{(t)},A_{i}^{(t)},V_{i}^{(t)}$
4.1.6 Tensorized Local Clustering Algorithm
Algorithm 2 Tensorized Federated Heat-Kernel Enhanced Multi-View Clustering (Local)
Input: Local tensor data $\mathcal{X}_{i}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{P}_{g},\mathbf{Q}_{g},\mathbf{R}_{g}\}$, personalization parameters $\{\lambda_{i},\rho_{i}\}$
Initialize:
$\mathcal{G}_{i}^{(0)}\leftarrow\lambda_{i}\mathcal{G}_{g}+(1-\lambda_{i})\mathcal{G}_{i}^{local}$
$\mathbf{P}_{i}^{(0)},\mathbf{Q}_{i}^{(0)},\mathbf{R}_{i}^{(0)}\leftarrow$ similarly personalized initialization
$\mathbf{v}_{i}^{(0)}\leftarrow\frac{1}{s_{i}}\mathbf{1}$
$t\leftarrow 0$
repeat
  $t\leftarrow t+1$
  // Compute tensorized heat-kernel coefficients
  Calculate $\delta_{j\ell h}^{(i)}$ using Eq. 61
  // Reconstruct cluster centers from decomposition
  $\mathcal{A}_{i}\leftarrow\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i}$
  // Update memberships
  Update $\mathcal{U}_{i}^{(t)}$ using Eq. 67
  // Update tensor factors using ALS
  Update $\mathcal{G}_{i}^{(t)}$ using Eq. 68
  Update $\mathbf{P}_{i}^{(t)}$ using Eq. 69
  Update $\mathbf{Q}_{i}^{(t)}$ using Eq. 70
  Update $\mathbf{R}_{i}^{(t)}$ using Eq. 71
  // Update view weights
  Update $\mathbf{v}_{i}^{(t)}$ using Eq. 72
  // Check convergence
  Compute $J_{i}^{tensor}$ using Eq. 63
until convergence or max iterations
return $\mathcal{U}_{i}^{(t)},\mathcal{G}_{i}^{(t)},\mathbf{P}_{i}^{(t)},\mathbf{Q}_{i}^{(t)},\mathbf{R}_{i}^{(t)},\mathbf{v}_{i}^{(t)}$
4.2 Federated Aggregation with Personalization
The federated learning framework coordinates local heat-kernel enhanced clustering results while preserving client personalization. The aggregation process operates on cluster centers and view importance weights.
4.2.1 Global Objective with Personalization
The global federated objective balances local clustering quality with personalization constraints:
$J_{global}=\sum_{i=1}^{M}\left[\omega_{i}J_{i}^{local}(U_{i},A_{i},V_{i})+\gamma\mathcal{R}_{person}(A_{i},A_{g})+\eta\mathcal{R}_{view}(V_{i},V_{g})\right]$ (73)
where:
  • $\omega_{i}$ represents client-specific weights based on data quality and quantity
  • $\mathcal{R}_{person}(A_{i},A_{g})=\sum_{h,k}\|A_{i}^{h}-A_{g}^{h}\|_{F}^{2}$ ensures coordination between local and global centers
  • $\mathcal{R}_{view}(V_{i},V_{g})=\|V_{i}-V_{g}\|_{2}^{2}$ maintains view weight consistency
  • $\gamma,\eta$ control personalization vs. coordination trade-offs
4.2.2 Privacy-Preserving Aggregation
To maintain privacy, clients share only aggregated statistics rather than raw cluster assignments:
Definition 1 (Federated Aggregation Protocol).
Each client $i$ computes and shares:
$S_{i}^{centers}$ $=\left\{\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}x_{j}^{h,(i)},\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}\right\}_{k,h}$ (74)
$S_{i}^{views}$ $=\left\{\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\text{KED}_{i}^{h}(x_{j}^{h,(i)},a_{k}^{h,(i)})\right\}_{h}$ (75)
$S_{i}^{quality}$ $=J_{i}^{local}\text{ and }n_{i}$ (76)
The server performs secure aggregation:
$A_{g}^{h,(t+1)}=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{centers,h}}{\sum_{i=1}^{M}\omega_{i}^{(t)}\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}}$ (77)
$V_{g}^{(t+1)}=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}V_{i}^{(t)}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (78)
where client weights are computed as:
$\omega_{i}^{(t)}=\frac{\exp(-\tau J_{i}^{local,(t)})}{\sum_{j=1}^{M}\exp(-\tau J_{j}^{local,(t)})}$ (79)
4.2.3 Personalized Update Mechanism
Clients receive global updates and perform personalized integration:
$A_{i}^{h,(t+1)}$ $=\lambda_{i}A_{i}^{h,(t)}+(1-\lambda_{i})A_{g}^{h,(t+1)}$ (80)
$V_{i}^{(t+1)}$ $=\rho_{i}V_{i}^{(t)}+(1-\rho_{i})V_{g}^{(t+1)}$ (81)
where $\lambda_{i},\rho_{i}\in[0,1]$ are client-specific personalization parameters that can be learned adaptively:
$\lambda_{i}^{(t+1)}=\lambda_{i}^{(t)}-\beta\nabla_{\lambda_{i}}J_{i}^{local}(\lambda_{i}A_{i}^{(t)}+(1-\lambda_{i})A_{g}^{(t)})$ (82)
4.3 Tensorized Federated Aggregation with Personalization
The federated learning framework coordinates local tensorized heat-kernel enhanced clustering results while preserving client personalization. The aggregation process operates on tensor decomposition factors: core tensors, factor matrices, and view importance weights.
4.3.1 Tensorized Global Objective with Personalization
The tensorized global federated objective balances local clustering quality with personalization constraints:
$J_{tensor\_global}=\sum_{i=1}^{M}\left[\omega_{i}J_{i}^{tensor}(\mathcal{U}_{i},\mathcal{G}_{i},\mathbf{P}_{i},\mathbf{Q}_{i},\mathbf{R}_{i},\mathbf{v}_{i})+\gamma\mathcal{R}_{tensor}(\mathcal{G}_{i},\mathbf{P}_{i},\mathbf{Q}_{i},\mathbf{R}_{i})+\eta\mathcal{R}_{view}(\mathbf{v}_{i},\mathbf{v}_{g})\right]$ (83)
where the tensorized regularization terms are:
$\mathcal{R}_{tensor}$ $=\|\mathcal{G}_{i}-\mathcal{G}_{g}\|_{F}^{2}+\|\mathbf{P}_{i}-\mathbf{P}_{g}\|_{F}^{2}+\|\mathbf{Q}_{i}-\mathbf{Q}_{g}\|_{F}^{2}+\|\mathbf{R}_{i}-\mathbf{R}_{g}\|_{F}^{2}$ (84)
$\mathcal{R}_{view}(\mathbf{v}_{i},\mathbf{v}_{g})$ $=\|\mathbf{v}_{i}-\mathbf{v}_{g}\|_{2}^{2}$ (85)
4.3.2 Privacy-Preserving Tensor Aggregation
To maintain privacy while enabling effective tensor factor coordination, clients share only aggregated tensor statistics rather than raw factors:
Definition 2 (Tensorized Federated Aggregation Protocol).
Each client $i$ computes and shares:
$S_{i}^{core}$ $=\left\{\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\mathcal{G}_{i,k,:,:}\right\}$ (86)
$S_{i}^{factors}$ $=\left\{\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}\mathbf{P}_{i,k,:},\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}\mathbf{Q}_{i,:,\ell},\sum_{j=1}^{n_{i}}(\mu_{jk}^{(i)})^{m}\mathbf{R}_{i,:,h}\right\}$ (87)
$S_{i}^{views}$ $=\left\{\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\text{TKED}_{i}(j,k,h)\right\}_{h}$ (88)
$S_{i}^{quality}$ $=J_{i}^{tensor}\text{ and }n_{i}$ (89)
The server performs secure tensor aggregation:
$\mathcal{G}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{core}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (90)
$\mathbf{P}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{factors,P}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (91)
$\mathbf{Q}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{factors,Q}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (92)
$\mathbf{R}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}S_{i}^{factors,R}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (93)
$\mathbf{v}_{g}^{(t+1)}$ $=\frac{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}\mathbf{v}_{i}^{(t)}}{\sum_{i=1}^{M}\omega_{i}^{(t)}n_{i}}$ (94)
4.3.3 Tensorized Personalized Update Mechanism
Clients receive global tensor updates and perform personalized integration across all tensor components:
$\mathcal{G}_{i}^{(t+1)}$ $=\lambda_{i}^{core}\mathcal{G}_{i}^{(t)}+(1-\lambda_{i}^{core})\mathcal{G}_{g}^{(t+1)}$ (95)
$\mathbf{P}_{i}^{(t+1)}$ $=\lambda_{i}^{P}\mathbf{P}_{i}^{(t)}+(1-\lambda_{i}^{P})\mathbf{P}_{g}^{(t+1)}$ (96)
$\mathbf{Q}_{i}^{(t+1)}$ $=\lambda_{i}^{Q}\mathbf{Q}_{i}^{(t)}+(1-\lambda_{i}^{Q})\mathbf{Q}_{g}^{(t+1)}$ (97)
$\mathbf{R}_{i}^{(t+1)}$ $=\lambda_{i}^{R}\mathbf{R}_{i}^{(t)}+(1-\lambda_{i}^{R})\mathbf{R}_{g}^{(t+1)}$ (98)
$\mathbf{v}_{i}^{(t+1)}$ $=\rho_{i}\mathbf{v}_{i}^{(t)}+(1-\rho_{i})\mathbf{v}_{g}^{(t+1)}$ (99)
where $\lambda_{i}^{core},\lambda_{i}^{P},\lambda_{i}^{Q},\lambda_{i}^{R},\rho_{i}\in[0,1]$ are component-specific personalization parameters.
4.4 Complete Tensorized Federated Algorithm
Algorithm 3 Personalized Federated Heat-Kernel Enhanced Tensorized Multi-View Clustering
Server Initialization:
 Initialize global tensor factors $\mathcal{G}_{g}^{(0)},\mathbf{P}_{g}^{(0)},\mathbf{Q}_{g}^{(0)},\mathbf{R}_{g}^{(0)}$ and view weights $\mathbf{v}_{g}^{(0)}$
 Set communication rounds $T$, local iterations $E$, tensor ranks $\{r_{2},r_{3}\}$
for $t=0,1,\ldots,T-1$ do
  Broadcast: Send $\mathcal{G}_{g}^{(t)},\mathbf{P}_{g}^{(t)},\mathbf{Q}_{g}^{(t)},\mathbf{R}_{g}^{(t)},\mathbf{v}_{g}^{(t)}$ to all clients
  Local Computation (in parallel for all clients):
  for each client $i\in\{1,\ldots,M\}$ do
   // Tensorize local multi-view data
   Organize $X_{i}=\{X_{i}^{h}\}_{h=1}^{s_{i}}$ into tensor $\mathcal{X}_{i}$
   // Personalized tensor factor initialization
   $\mathcal{G}_{i}^{(0)}\leftarrow\lambda_{i}^{core}\mathcal{G}_{g}^{(t)}+(1-\lambda_{i}^{core})\mathcal{G}_{i}^{local}$
   $\mathbf{P}_{i}^{(0)}\leftarrow\lambda_{i}^{P}\mathbf{P}_{g}^{(t)}+(1-\lambda_{i}^{P})\mathbf{P}_{i}^{local}$
   $\mathbf{Q}_{i}^{(0)}\leftarrow\lambda_{i}^{Q}\mathbf{Q}_{g}^{(t)}+(1-\lambda_{i}^{Q})\mathbf{Q}_{i}^{local}$
   $\mathbf{R}_{i}^{(0)}\leftarrow\lambda_{i}^{R}\mathbf{R}_{g}^{(t)}+(1-\lambda_{i}^{R})\mathbf{R}_{i}^{local}$
   $\mathbf{v}_{i}^{(0)}\leftarrow\rho_{i}\mathbf{v}_{g}^{(t)}+(1-\rho_{i})\mathbf{v}_{i}^{local}$
   // Local tensorized heat-kernel enhanced clustering
   for $e=1,2,\ldots,E$ do
    Compute $\delta_{j\ell h}^{(i)}$ using Eq. 61
    Reconstruct $\mathcal{A}_{i}\leftarrow\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i}$
    Update $\mathcal{U}_{i}$ using Eq. 67
    Update $\mathcal{G}_{i}$ using Eq. 68
    Update $\mathbf{P}_{i}$ using Eq. 69
    Update $\mathbf{Q}_{i}$ using Eq. 70
    Update $\mathbf{R}_{i}$ using Eq. 71
    Update $\mathbf{v}_{i}$ using Eq. 72
   end for
   // Compute tensorized aggregation statistics
   Compute $S_{i}^{core},S_{i}^{factors},S_{i}^{views},S_{i}^{quality}$
   Send statistics to server
  end for
  Server Tensor Aggregation:
  Compute client weights $\omega_{i}$ using tensor quality measures
  Update $\mathcal{G}_{g}^{(t+1)}$ using Eq. 90
  Update $\mathbf{P}_{g}^{(t+1)}$ using Eq. 91
  Update $\mathbf{Q}_{g}^{(t+1)}$ using Eq. 92
  Update $\mathbf{R}_{g}^{(t+1)}$ using Eq. 93
  Update $\mathbf{v}_{g}^{(t+1)}$ using Eq. 94
  Adaptive Personalization:
  Update tensor-specific personalization parameters $\{\lambda_{i}^{core},\lambda_{i}^{P},\lambda_{i}^{Q},\lambda_{i}^{R},\rho_{i}\}$
end for
return Global tensor model $(\mathcal{G}_{g}^{(T)},\mathbf{P}_{g}^{(T)},\mathbf{Q}_{g}^{(T)},\mathbf{R}_{g}^{(T)},\mathbf{v}_{g}^{(T)})$ and personalized models $\{(\mathcal{G}_{i}^{(T)},\mathbf{P}_{i}^{(T)},\mathbf{Q}_{i}^{(T)},\mathbf{R}_{i}^{(T)},\mathbf{v}_{i}^{(T)})\}_{i=1}^{M}$
5 Theoretical Analysis
5.1 Convergence Properties
Theorem 3 (Local Convergence).
For each client $i$, the local heat-kernel enhanced multi-view fuzzy clustering algorithm converges to a local minimum of the objective function $J_{i}^{local}$.
Proof.
The proof follows from the monotonic decrease property of alternating optimization. Each update step:
  1. 1.
    Membership updates (Eq. 26) minimize $J_{i}^{local}$ w.r.t. $U_{i}$ for fixed $A_{i},V_{i}$
  2. 2.
    Center updates (Eq. 27) minimize $J_{i}^{local}$ w.r.t. $A_{i}$ for fixed $U_{i},V_{i}$
  3. 3.
    View weight updates (Eq. 28) minimize $J_{i}^{local}$ w.r.t. $V_{i}$ for fixed $U_{i},A_{i}$
Since $J_{i}^{local}$ is bounded below by zero and decreases monotonically, convergence is guaranteed. ∎
Theorem 4 (Global Convergence).
Under mild regularity conditions, the federated algorithm converges to a stationary point of the global objective function $J_{global}$.
Proof.
The global convergence follows from the convergence theory of federated optimization. The key insight is that each local update corresponds to a proximal gradient step on the global objective with personalization regularization, ensuring convergence to a stationary point. ∎
5.2 Privacy Guarantees
Theorem 5 (Differential Privacy).
The federated aggregation protocol provides $(\epsilon,\delta)$-differential privacy when clients add calibrated noise to their shared statistics.
Proof.
The shared statistics $S_{i}^{centers},S_{i}^{views},S_{i}^{quality}$ have bounded sensitivity. Adding Gaussian noise with variance $\sigma^{2}=\frac{2\Delta^{2}\log(1.25/\delta)}{\epsilon^{2}}$ where $\Delta$ is the sensitivity, ensures $(\epsilon,\delta)$-differential privacy. ∎
Proposition 1 (Communication Efficiency).
The federated protocol requires $\mathcal{O}(c\cdot\bar{d}\cdot\bar{s})$ communication per round, where $\bar{d}$ and $\bar{s}$ are average feature dimensions and views across clients.
5.3 Computational Complexity
Theorem 6 (Time Complexity).
The local clustering complexity per client $i$ per iteration is $\mathcal{O}(n_{i}cs_{i}D_{i})$ where $D_{i}=\sum_{h=1}^{s_{i}}d_{h}^{(i)}$ is the total local dimensionality.
The federated approach maintains the same asymptotic complexity as centralized multi-view clustering while providing personalization and privacy benefits.
5.4 Tensorized Complexity Analysis
Theorem 7 (Tensorized Time Complexity).
The tensorized local clustering complexity per client $i$ per iteration is $\mathcal{O}(n_{i}cr_{2}r_{3}+n_{i}cD_{i}s_{i})$ where $r_{2},r_{3}$ are the tensor decomposition ranks.
Proof.
The complexity breakdown includes:
  1. 1.
    Heat-kernel coefficient computation: $\mathcal{O}(n_{i}D_{i}s_{i})$
  2. 2.
    Tensor reconstruction: $\mathcal{O}(cr_{2}r_{3}+cD_{i}r_{2}+cs_{i}r_{3})$
  3. 3.
    Membership updates: $\mathcal{O}(n_{i}cs_{i}D_{i})$
  4. 4.
    Tucker decomposition updates: $\mathcal{O}(n_{i}cr_{2}r_{3})$ per factor using ALS
The dominant term is $\mathcal{O}(n_{i}cD_{i}s_{i})$ when $r_{2},r_{3}\ll\min(D_{i},s_{i})$. ∎
Theorem 8 (Tensorized Communication Efficiency).
The tensorized federated protocol requires $\mathcal{O}(cr_{2}r_{3}+c^{2}+\bar{D}r_{2}+\bar{s}r_{3})$ communication per round, compared to $\mathcal{O}(c\bar{D}\bar{s})$ for the non-tensorized version.
The tensorized approach provides significant communication savings when $r_{2}\ll\bar{D}$ and $r_{3}\ll\bar{s}$, which is typically the case in practice.
Lemma 1 (Tensorized Approximation Quality).
The reconstruction error of the tensorized cluster centers is bounded by:
$\|\mathcal{A}_{i}-\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i}\|_{F}^{2}\leq\epsilon_{tensor}$ (100)
where $\epsilon_{tensor}$ depends on the chosen tensor ranks and the intrinsic tensor rank of the true cluster center tensor.
5.5 Personalization-Accuracy Trade-off
Lemma 2 (Personalization Bound).
The personalization error is bounded by:
$\|A_{i}-A_{g}\|_{F}^{2}\leq\frac{\lambda_{i}^{2}}{1-\lambda_{i}^{2}}\|A_{i}^{local}-A_{g}\|_{F}^{2}$ (101)
where $A_{i}^{local}$ is the purely local optimum.
This bound shows that personalization parameters $\lambda_{i}$ control the trade-off between local specialization and global coordination.
6 Conclusion
This paper presents a novel personalized federated learning framework that integrates heat-kernel enhanced multi-view clustering with tensor-based global coordination. The key contributions include:
  1. 1.
    Heat-kernel Enhanced Federated Clustering: We adapt heat-kernel coefficients to federated multi-view clustering, enabling more effective capture of complex local patterns while maintaining global coordination.
  2. 2.
    Personalized Aggregation Mechanism: Our framework provides adaptive personalization that balances local specialization with global knowledge sharing through learnable personalization parameters.
  3. 3.
    Privacy-Preserving Protocol: The aggregation protocol ensures differential privacy while maintaining clustering effectiveness through carefully designed statistical sharing.
  4. 4.
    Comprehensive Theoretical Analysis: We provide convergence guarantees, privacy bounds, and complexity analysis that establish the theoretical foundation of the approach.
6.1 Future Directions
Several promising research directions emerge from this work:
  • Dynamic View Discovery: Extending the framework to handle dynamically changing view structures across federated rounds.
  • Hierarchical Federated Clustering: Investigating multi-level federated architectures for large-scale deployment.
  • Adversarial Robustness: Analyzing and improving robustness against malicious clients and data poisoning attacks.
  • Continuous Learning: Adapting the framework for federated continual learning scenarios where data distributions evolve over time.
  • Cross-modal Applications: Exploring applications to federated learning with fundamentally different data modalities (text, images, sensors, etc.).
The proposed framework opens new possibilities for privacy-preserving, personalized machine learning on distributed multi-view data, with broad applications in healthcare, IoT, and collaborative intelligence systems.