### v2

# Personalized Enhanced Federated Multi-View Clustering via Heat-Kernel Tensor Decomposition

Abstract
This paper proposes a personalized federated learning framework integrating heat-kernel enhanced tensorized multi-view fuzzy c-means clustering with tensor decomposition techniques. The approach combines heat-kernel coefficients adapted from quantum field theory with PARAFAC2 and Tucker decomposition to transform distance metrics and efficiently represent high-dimensional multi-view structures. Two main algorithms, FedHK-PARAFAC2 and FedHK-Tucker, are developed to extract shared and view-specific features while preserving inter-view relationships. The framework addresses data heterogeneity, privacy preservation, and communication efficiency challenges in federated learning environments. Theoretical analysis provides convergence guarantees, privacy bounds, and complexity analysis. The integration of heat-kernel methods with tensor decomposition in a federated setting offers a novel approach for effective multi-view data analysis while ensuring data privacy.
Keywords: Federated learning, Multi-view clustering, Heat-kernel methods, Tensor decomposition, PARAFAC2, Tucker decomposition, Privacy preservation.
1 Introduction
Large-scale data centers become increasingly important as the Internet of Things (IoT) continues to grow, with more devices being massively connected to the internet than ever before. This phenomenal transition of data to distributed environments affects the need for investigating optimal approaches to process heterogeneous information from multiple devices while ensuring security and efficient data exchange. Conventional unsupervised techniques cannot effectively handle these higher-order or higher-dimensional data structures inherent in modern federated systems where distributed clients possess different data modalities, feature dimensions, and local patterns.
Tensor-based machine learning approaches have emerged as powerful tools for analyzing multi-dimensional data structures, enabling the capture of complex relationships and interactions across multiple modes. Tensors, as generalizations of matrices to higher dimensions, provide a natural framework for representing multi-view data, where each view corresponds to a different modality or feature set. For instance, in image processing, a color image can be represented as a third-order tensor with dimensions corresponding to height, width, and color channels (Red, Green, Blue) as illustrated in Figure 1.
Figure 1: The Illustration of A 3-Dimensional of a Red-Green-Blue (RGB) Image
As can be seen in Figure 1, a color image is represented as a 3-dimensional tensor, where each slice along the third dimension corresponds to one of the RGB color channels. This tensor representation allows for the preservation of spatial and color information, enabling more effective analysis and processing of image data. RGB image is stored in a 3-Dimensional array as $\text{channels}\times\text{height}\times\text{width}$ or $\text{height}\times\text{width}\times\text{channels}$. The $7\times 7\times 3$ tensor in Figure 1 represents a small RGB image with a height and width of 7 pixels and 3 color channels. Each element in the tensor corresponds to the intensity value of a specific color channel at a particular pixel location.
In federated learning environments, clients often possess multi-view (MV) data that can be naturally represented as tensors. For example, in healthcare applications, different hospitals may collect various types of patient data, such as medical images, electronic health records, and genomic data, each representing a different view of the patient’s health status. By leveraging tensor-based representations, federated learning algorithms can effectively capture the multi-modal nature of the data while preserving privacy and minimizing communication overhead. Singular value decomposition (SVD) and principal component analysis (PCA) [welke2006fast, sinaga2021entropy] are classical matrix factorization techniques that have been widely used for dimensionality reduction and feature extraction. However, these methods are limited to two-dimensional data structures and cannot fully exploit the multi-way relationships present in tensor data. Tensor decomposition methods, such as CANDECOMP/PARAFAC (CP) [harshman1970foundations] and Tucker decomposition [tucker1966some], extend these concepts to higher-order tensors, enabling more comprehensive analysis of multi-view data.
From the perspective of MV clustering, each component of feature-view stored as a two-order tensor (matrix) can be naturally represented as a three-order tensor by stacking the matrices along a new dimension corresponding to the views. This tensor representation allows for the preservation of inter-view relationships and facilitates the application of tensor decomposition techniques to extract shared and view-specific features. By leveraging tensor algebra, MV clustering algorithms can effectively capture the complex interactions between different views, leading to improved clustering performance and robustness.
The analogy of diversity and heterogeneity on multiple representations of the same information is precise [giunchiglia2022representation] . In this sense, multi-view data can be viewed as different perspectives or modalities of the same underlying phenomenon, each providing unique insights and complementary information. For example, in social network analysis, user profiles may include text data from posts, images shared, and interaction patterns with other users. Each of these views captures different aspects of user behavior and preferences, contributing to a more holistic understanding of the social dynamics.
This paper presents two novel algorithms, namely FedHK-PARAFAC2 and FedHK-Tucker, that integrate heat-kernel enhanced tensorized multi-view fuzzy c-means clustering with advanced tensor decomposition techniques in a federated learning framework. The proposed methods leverage heat-kernel coefficients adapted from quantum field theory to transform conventional distance metrics, enabling the capture of local geometric relationships within multi-view data distributed across heterogeneous clients. By incorporating PARAFAC2 and Tucker decomposition, the algorithms efficiently represent high-dimensional multi-view structures while addressing challenges of data heterogeneity, privacy preservation, and communication efficiency inherent in federated learning environments.
2 Related Work
In order to provide a comprehensive overview of the contributions of this study, a review of the extant literature was conducted across four key domains: federated learning and clustering, multi-view clustering, heat kernel methods and quantum-inspired learning, and tensor decomposition techniques. This review underscores foundational works and recent advancements in the field, while concurrently identifying gaps that the proposed framework aims to address.
2.1 Federated Learning and Clustering
Pattern recognition and clustering in federated learning (FL) environments have garnered significant attention due to the increasing need for privacy-preserving distributed data analysis. Federated learning, introduced by McMahan et al. [mcmahan2017communication], enables collaborative model training across decentralized clients without sharing raw data. While initial research focused on supervised learning tasks, recent efforts have extended federated learning to unsupervised clustering scenarios. Early work by Dennis et al. [dennis2021heterogeneity] explored federated k-means clustering, demonstrating the feasibility of distributed clustering with local updates and global aggregation. Ghosh et al. [ghosh2002consensus] proposed consensus-based approaches for federated hierarchical clustering, addressing challenges of data heterogeneity and communication efficiency. Makhija et al. [makhija2022federated] introduced federated fuzzy c-means clustering, leveraging soft assignments to enhance robustness against noise and outliers. However, these methods often assume homogeneous data distributions and do not fully exploit multi-view data structures prevalent in modern applications. Recent advancements have focused on personalized federated clustering to accommodate client heterogeneity. Caldarola et al. [caldarola2021cluster] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. [li2022hpfl] developed meta-learning approaches for personalized cluster initialization. Privacy-preserving aspects have been addressed through differential privacy mechanisms [kairouz2021advances] and secure aggregation protocols [bonawitz2017practical], ensuring that sensitive information remains protected during the clustering process. Communication efficiency has been tackled through gradient compression [konecny2016federated], sparsification techniques [aji2017sparse], and model quantization [reisizadeh2020fedpaq], reducing the overhead associated with transmitting model updates across clients.
2.2 Multi-View Clustering
Multi-view clustering (MVC) has emerged as a powerful paradigm for leveraging complementary information from multiple data representations to enhance clustering performance. Early work by Bickel and Scheffer [bickel2004multi] established the theoretical foundations for multi-view learning, demonstrating that integrating multiple views can lead to improved clustering accuracy. Classical MVC approaches can be categorized into co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as those proposed by Kumar et al. [kumar2011co], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [hotelling1992relations] and shared subspace clustering [wang2019multiple], project multiple views into a common representation space for unified clustering. Graph-based MVC has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. [wang2019gmc] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. [nie2016parameter] developed parameter-free auto-weighted multiple graph learning. Deep learning approaches have revolutionized MVC through end-to-end representation learning. Zhang et al. [zhang2019ae2] introduced auto-encoder based multi-view clustering, while Zhu et al. [zhu2019multi] developed deep multi-view clustering networks with view correlation discovery. Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. [jiang2014collaborative] and Wang et al. [wang2017multi], introducing weighted aggregation schemes and view-specific fuzzification parameters. Sinaga et al. [sinaga2025parameter] proposed a parameter-free multi-view fuzzy clustering algorithm that automatically learns view weights and fuzzification parameters, enhancing clustering robustness and interpretability. However, these methods typically require centralized data access and are not designed for federated environments.
The federated multi-view clustering is first introduced by Chen et al. [chen2022federated] for federated multi-view representation learning and Liu et al. [liu2023federated] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication. Yang et al. [yang2024federated] proposed a federated multi-view clustering framework that combines local clustering with global consensus, but it lacks the integration of heat-kernel methods and advanced tensor decomposition techniques. Earlier, Sinaga [sinaga2025fedhk] proposed a heat-kernel enhanced federated multi-view clustering framework, but it did not incorporate tensor decomposition methods for efficient multi-view representation. Thus, there is a need for advanced federated multi-view clustering approaches that leverage tensor decomposition and heat-kernel methods to address data heterogeneity, privacy preservation, and communication efficiency challenges in distributed settings.
2.3 Heat Kernel Methods and Quantum-Inspired Learning
Heat kernel methods have emerged as powerful tools for capturing local geometric structure in high-dimensional data spaces. Originally developed in differential geometry and quantum field theory, heat kernels provide a natural way to encode multi-scale geometric information through diffusion processes on manifolds. The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. [ham2004kernel] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon [coifman2006diffusion] established theoretical foundations for diffusion-based data analysis. In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona [zelnik2004self] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg [von2007tutorial] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. [havlivcek2019supervised], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [carleo2017solving] and quantum neural networks [biamonte2017quantum]. The integration of heat kernel coefficients with clustering algorithms represents a natural extension of kernel methods to soft clustering scenarios. By incorporating heat diffusion properties, clustering algorithms can better capture local data manifold structure and improve robustness to noise and outliers.
2.4 Tensor Decomposition Methods
Tensor decomposition has emerged as a powerful paradigm for analyzing high-dimensional multi-modal data, providing compact representations that capture multilinear relationships inherent in complex datasets. The field has evolved from classical matrix factorization techniques to sophisticated tensor algebra frameworks capable of handling arbitrary N-way data structures. Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang [carroll1970analysis] and Harshman [harshman1970foundations], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. Tucker decomposition, proposed by Tucker [tucker1966some], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [de2000multilinear] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems. Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [liu2012tensor] and Yuan and Zhang [yuan2016tensor], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [huang2014robust] provide resilience against outliers and corrupted entries. Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [lee2000algorithms] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [shashua2005non], enabling decompositions with non-negativity constraints that often yield more interpretable factors. The integration of tensor decomposition with machine learning has led to tensorized neural networks [novikov2015tensorizing], tensor-based kernel methods [signoretto2011kernel], and multi-task learning frameworks [wimalawarne2014multitask]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power. In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Zhou et al. [cichocki2015tensor] developed tensor-based multiway arrays for bid data sets, while Benjamin et al. [benjamin2024tensor] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.
3 Mathematical Preliminaries
This section presents a thorough mathematical formulation of the fundamental concepts and notations employed in this study. The section commences with a review of the principles of federated learning and clustering. This is followed by an examination of multi-view clustering techniques. Subsequently, the section explores heat kernel methods and their applications in quantum-inspired learning. Finally, the discussion will turn to tensor decomposition methods relevant to the proposed framework.
Notation and Mathematical Framework: Let $X=\{x_{1},\ldots,x_{n}\}$ be the collection of $n$ data points in $\mathbb{R}^{d}$ with $x_{i}=[x_{ij}]_{n\times d}$, $i=1,\ldots,n$, $j=1,\ldots,d$; $A=\{a_{1},\ldots,a_{c}\}$ be the set of cluster centers with $a_{k}=[a_{kj}]_{c\times d}$ and $k=1,\ldots,c$, and $U=\{\mu_{1},\ldots,\mu_{c}\}$ be the cluster membership matrix with $\mu_{k}=[\mu_{ik}]_{n\times c}$.
In the multi-view scenario, $h$ denotes the view index, $h=1,\ldots,s$, and the feature dimensionality of the $h$-th view is denoted as $d_{h}$. Thus, $X^{h}=\{x_{1}^{h},\ldots,x_{n}^{h}\}$ defines the data input $X$ in the $h$-th view with $x_{i}^{h}=[x_{ij}^{h}]_{n\times d_{h}}$. Hence, $A^{h}=\{a_{1}^{h},\ldots,a_{c}^{h}\}_{h=1}^{s}$ represents the cluster centers of the $h$-th view with $a_{k}^{h}=[a_{kj}^{h}]_{c\times d_{h}}$.
3.1 Federated Learning and Clustering
Federated learning, pioneered by McMahan et al. [mcmahan2017communication], enables collaborative machine learning across distributed clients without centralizing raw data. The seminal FedAvg algorithm established the foundation for aggregating local model updates through weighted averaging. However, traditional federated learning primarily focuses on supervised tasks, with federated clustering receiving comparatively less attention despite its practical importance in distributed data analytics.
Konečnỳ et al. [konevcny2016federated] introduced federated optimization as a method to train a centralized model using decentralized data stored across numerous devices, emphasizing communication efficiency due to limited connectivity and privacy concerns. They proposed modifications to existing algorithms to adapt them to federated optimization settings. Li et al. [li2020acceleration] adopted the foundational ideas of federated optimization and extended them by integrating advanced techniques like gradient compression, acceleration, and variance reduction to address communication bottlenecks and improve the scalability of federated learning systems
Early work on federated clustering by Dennis et al. [dennis2021heterogeneity], Ghosh et al. [ghosh2002consensus], and Makhija et al. [makhija2022federated] demonstrated the feasibility of distributed k-means and hierarchical clustering. These approaches typically involve local clustering followed by global aggregation of cluster centers or assignments. However, these methods assume homogeneous data distributions and fail to address the inherent heterogeneity in federated environments.
Recent advances have introduced personalized federated clustering to address client heterogeneity. Caldarola et al. [caldarola2021cluster] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. [li2022hpfl] developed meta-learning approaches for personalized cluster initialization. These works, however, are limited to single-view data and do not leverage the rich multi-modal information available in modern federated systems.
Privacy-preserving aspects of federated clustering have been addressed through differential privacy mechanisms [kairouz2021advances] and secure aggregation protocols [bonawitz2017practical]. Geyer et al. [geyer2017differentially] demonstrated how differential privacy can be integrated into federated clustering while maintaining acceptable utility. However, these privacy mechanisms often come at the cost of clustering accuracy, particularly in high-dimensional settings.
The communication efficiency challenge in federated clustering has been tackled through gradient compression [konecny2016federated], sparsification techniques [aji2017sparse], and model quantization [reisizadeh2020fedpaq]. While these methods reduce communication overhead, they typically do not exploit the inherent structure of multi-view tensor data for more efficient representation and transmission.
3.2 Multi-View Clustering
Multi-view clustering leverages complementary information from multiple data representations to improve clustering performance. Early work by Bickel and Scheffer [bickel2004multi] established the theoretical foundations for multi-view learning, demonstrating superior performance when views provide complementary information about the underlying data structure.
Classical multi-view clustering approaches can be categorized into three main paradigms: co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as the work by Kumar et al. [kumar2011co], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [hotelling1992relations] and shared subspace clustering [wang2019multiple], project multiple views into a common representation space for unified clustering.
Graph-based multi-view clustering has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. [wang2019gmc] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. [nie2016parameter] developed parameter-free auto-weighted multiple graph learning. These methods construct view-specific graphs and learn optimal view weights for clustering.
Deep learning approaches have revolutionized multi-view clustering through end-to-end representation learning. Zhang et al. [zhang2019ae2] introduced auto-encoder based multi-view clustering, while Zhu et al. [zhu2019multi] developed deep multi-view clustering networks with view correlation discovery. However, these methods typically require centralized data access and are not designed for federated environments.
Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. [jiang2014collaborative] and Wang et al. [wang2017multi], introducing weighted aggregation schemes and view-specific fuzzification parameters. These soft clustering approaches provide uncertainty quantification but lack the scalability and privacy considerations required for federated deployment.
The mathematical foundations of multi-view clustering build upon single-view clustering formulations. In this study, we focus on soft clustering approaches that provide probabilistic cluster assignments, enabling better uncertainty quantification in federated environments. The objective function of standard fuzzy c-means (FCM) or soft clustering can be expressed as:
$J_{FCM}(U,A)=\sum_{i=1}^{n}\sum_{k=1}^{c}\mu_{ik}^{m}\|x_{i}-a_{k}\|^{2}$ (1)
where $U$ is the membership matrix of data sets $X$ in clusters $c$, AA is the cluster centers matrices of the $j$-th feature in clusters $c$, and $m$ is the fuzzifier to control the distribution of membership matrix $U$. Eq. 1 is designed to address data with a single representation, which is not compatible to solve a problem with multiple representations. To tackle this, $s$ number of views is introduced, such as indicator $h=1,\ldots,s$ enabled the representation of data X from $s$ views. In this way, the clustering problems based on multiple representation data $X^{h}$ can be solved by using the following equation:
$J_{MVC}(U^{h},A^{h})=\sum_{h=1}^{s}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{h})^{m}\|x_{i}^{h}-a_{k}^{h}\|^{2}$ (2)
As displayed in Eq. 2, the memberships of the $h$-th view data in $c$ clusters $\mu_{ik}^{h}$ is computed locally. Usually, to get a global memberships of data view $s$, we have an additional step called as aggregation of collected memberships across data views $s$. There are multiple ways to aggregate the memberships of $s$ data views, i.e., using the averagation approach or taking a product of memberships of the $i$-th data point across all views $s$ in clusters $c$ and divide it with the number of data points. If we consider a weight factor $v_{h}$, we can multiply each weight factors with the memberships and divide it by the number of data views, etc. As we focused on a client’ multi-view data, we simplify the computation for each view memberships by directly implementing the global memberships from an early iteration. In this way, we can modify Eq. 2 in the following way
$J_{{MVC}_{2}}(U^{*},A^{h})=\sum_{h=1}^{s}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}\|x_{i}^{h}-a_{k}^{h}\|^{2}$ (3)
The objective is to provide a comprehensive mathematical objective function that can effectively handle multi-view data distributed across multiple clients using tensor algebra, taking into account non-linearity compatibilities. Consequently, the distance between data points and cluster centers across data views in Euclidean space is replaced by a kernel Euclidean distance (KED). The following definition is proposed for the mathematical expression of KED:
$\text{KED}(x_{ij}^{h},a_{kj}^{h})=\bigg\{1-\text{exp}\bigg(-\sum_{j=1}^{d_{h}}\delta_{ij}^{h}(x_{ij}^{h}-a_{kj}^{h})^{2}\bigg)\bigg\}$ (4)
where $\delta_{ij}^{h}$ is a heat kernel coefficient (HKC) to control the Euclidean distance between data points $x_{ij}^{h}$ and cluster centers $a_{kj}^{h}$ across data views in kernel space. In this way, we can modify Eq. 3 in the following way
$J_{KMVC}(U^{*},A^{h})=\sum_{h=1}^{s}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}\text{KED}(x_{ij}^{h},a_{kj}^{h})$ (5)
subject to:
$\sum_{k=1}^{c}\mu_{ik}^{*}=1,\mu_{ik}^{*}\in[0,1]$
As illustrated in Eq. 5, it is imperative to prioritize all data views equally. This approach is not conducive to the efficient resolution of multiple resources data. Given that divergent perspectives correspond to disparate manifestations of feature behavior, it becomes imperative to allocate a distinct weight, denoted as $v_{h}$, to each data view, denoted as $x_{i}^{h}$. It is imperative to underscore that this allocation must be completed, in accordance with the principle that $\sum_{h=1}^{s}v_{h}=1$. By incorporating the weight factors, denoted by $v_{h}$, into the analysis, a modification to the Eq. 5 can be formulated as follows:
$J_{E-KMVC}(V_{h},U^{*},A^{h})=\sum_{h=1}^{s}v_{h}^{\alpha}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}\bigg\{1-\text{exp}\big(-\sum_{j=1}^{d_{h}}\delta_{ij}^{h}(x_{ij}^{h}-a_{kj}^{h})^{2}\big)\bigg\}$ (6)
subject to:
$\sum_{k=1}^{c}\mu_{ik}^{*}=1,\mu_{ik}^{*}\in[0,1],\quad\sum_{h=1}^{s}v_{h}=1,v_{h}\in[0,1]$
where $\alpha$ is an exponent parameter to control the distribution of data views during the clustering processes.
3.3 Heat Kernel Methods and Quantum-Inspired Learning
Heat kernel methods have emerged as powerful tools for capturing local geometric structure in high-dimensional data spaces. Originally developed in differential geometry and quantum field theory, heat kernels provide a natural way to encode multi-scale geometric information through diffusion processes on manifolds.
The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. [ham2004kernel] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon [coifman2006diffusion] established theoretical foundations for diffusion-based data analysis.
In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona [zelnik2004self] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg [von2007tutorial] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. [havlivcek2019supervised], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [carleo2017solving] and quantum neural networks [biamonte2017quantum].
The integration of heat kernel coefficients with clustering algorithms represents a natural extension of kernel methods to soft clustering scenarios. By incorporating heat diffusion properties, clustering algorithms can better capture local data manifold structure and improve robustness to noise and outliers.
3.4 Federated Multi-View Learning and Privacy Preservation
The intersection of federated learning and multi-view clustering presents unique challenges that are not adequately addressed by existing literature. Most federated learning approaches assume single-view data and homogeneous client populations, while multi-view clustering methods typically require centralized data access.
Recent work on federated multi-view learning includes approaches by Chen et al. [chen2022federated] for federated multi-view representation learning and Liu et al. [liu2023federated] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication.
Privacy preservation in federated multi-view settings requires careful consideration of information leakage across views and clients. Differential privacy mechanisms must be adapted to handle the multi-view tensor structure while maintaining clustering utility. Secure aggregation protocols need to be extended to handle tensor factor aggregation efficiently.
The communication efficiency challenge is particularly acute in federated multi-view scenarios due to the high-dimensional nature of multi-view tensor data. Existing compression techniques are not optimized for tensor structures and may destroy important multilinear relationships. This motivates the need for tensor-aware compression and aggregation schemes.
The subsequent discourse will encompass the tensorization and activation of equation Eq. 6. In the following sections, the open challenges will be addressed by means of a proposed heat-kernel-enhanced tensorized framework in federated environments.
3.5 Tensor Decomposition Methods
Tensor decomposition has emerged as a powerful paradigm for analyzing high-dimensional multi-modal data, providing compact representations that capture multilinear relationships inherent in complex datasets. The field has evolved from classical matrix factorization techniques to sophisticated tensor algebra frameworks capable of handling arbitrary N-way data structures.
Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang [carroll1970analysis] and Harshman [harshman1970foundations], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. The uniqueness property of CPD makes it particularly valuable for component analysis and factor identification in multi-way data.
Tucker decomposition, proposed by Tucker [tucker1966some], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [de2000multilinear] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems.
Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [liu2012tensor] and Yuan and Zhang [yuan2016tensor], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [huang2014robust] provide resilience against outliers and corrupted entries.
Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [lee2000algorithms] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [shashua2005non], enabling decompositions with non-negativity constraints that often yield more interpretable factors.
The integration of tensor decomposition with machine learning has led to tensorized neural networks [novikov2015tensorizing], tensor-based kernel methods [signoretto2011kernel], and multi-task learning frameworks [wimalawarne2014multitask]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power.
In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Zhou et al. [cichocki2015tensor] developed tensor-based multiway arrays for bid data sets, while Benjamin et al. [benjamin2024tensor] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.
A tensor is a higher-order array that represents multi-dimensional data structures. Scalar is a zero-order tensor; Vector is a 1-dimensional array, also known as tensor of order one and denoted by boldface lowercase letter, e.g., $\mathbf{x}$; A matrix $\mathbf{X}$ is a 2-way array or tensor of order two with two indices $(m,n)$, where $m$ indicates the row dimension and $n$ indicates the column dimension; While a tensor $\mathcal{X}$ is a generalization of matrices with three or more indices $(i,j,k,\ldots)$, considered as 3-way arrays or higher-way or N-way arrays (higher-order) tensors.
When $X\in\mathbb{R}^{n\times d_{h}\times s}$ represents multi-view data in tensor space, the notation $\mathcal{X}_{ijh}$ denotes the score or element of the $i$-th data point on the $j$-th feature component at the $h$-th view. Figure 2 illustrates the representation of multi-view data in tensor space. As clearly shown, multi-view data can be viewed as a 3-dimensional tensor $\mathcal{X}=\{x_{ijh}\}_{i=1,j=1,h=1}^{n,d_{h},s}$ or equivalently as a set of 2D matrices $\mathcal{X}=\{X^{1},\ldots,X^{s}\}$ with size $n\times d_{h}$.
Figure 2: The Illustration of MV Data in Tensor Space
Tensor Operations and Decomposition Preliminaries: The elements of $N$-way arrays can be reordered into matrices or vectors through matricization and vectorization operations. The matricization is reordering tensor’ elements into a matrix, while vectorization is reordering tensor’ elements into a vector. Canonical polyadic decomposition (CPD) or parallel factors (PARAFAC) [harshman1994parafac, carroll1970analysis], Tucker decomposition (TD), and non-negative tensor factorization (NTF) are well-known approaches to reorder a tensor into a matrix or vector. PARAFAC or CPD holds multilinearity assumptions such as handle a sum of rank-one tensors by targeting the same lengths. In this sense, the original PARAFAC decomposition can reveal uncorrelated (orthogonal) and correlated (oblique) factors in system-variation data matrices, but failed to describe correlated factors when applied to the cross-product matrices computed from one data [harshman1972parafac2, kiers1999parafac2]. PARAFAC2, on the other hand has the advantageous to handle irregular tensors by relaxing the canonical polyadic (CP) model [harshman1972parafac2]. PARAFAC2 enables the three-mode model, often implemented into subspace clustering algorithms taking advantageous of correlations between pair of factors. The PARAFAC2 normalize the generalization of PARAFAC to deal both with uncorrelated and correlated factors.
3.5.1 The Basic Background of Tensors
This subsection provides some definitions and concepts of tensor algebra, most of overviews and notations are borrowed from Kolda et al. [kolda2009tensor].
Definition 1: The $N$-th order of tensor $\mathcal{A}$ is expressed as
$\mathcal{A}=[\mathcal{A}_{g_{1},g_{2},\ldots,g_{n}}]\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ (7)
where the $(g_{1},g_{2},\ldots,g_{n})$-th elements of the tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is containing $(\mathcal{A})_{g_{1},g_{2},\ldots,g_{n}}$ or $a_{g_{1},g_{2},\ldots,g_{n}}$ for indices $g_{1}=\{1,2,\ldots,G_{1}\}$, $g_{2}=\{1,2,\ldots,G_{2}\}$, and $g_{n}=\{1,2,\ldots,G_{N}\}$.
Definition 2: The outer product of $N$ vectors of rank-one $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is defined as
$\mathcal{A}=a^{(1)}\circ a^{(2)}\circ\ldots\circ a^{(N)}$ (8)
where $\circ$” represents the vector outer product. Fig. 3 presented the illustration of rank-one third-order tensor $\mathcal{A}=x\circ y\circ z$. The $(i,j,k)$ element of $\mathcal{A}$ is given by $a_{ijk}=x_{i}y_{j}z_{k}$. In terms of element wise, the 3-way tensor can be written as $a_{ijk}\approx\sum_{r=1}^{R}x_{ir}y_{jr}z_{kr}$.
Figure 3: The Illustration of Rank-One Third-Order Tensor $\mathcal{A}=x\circ y\circ z$
Definition 3: The $n$-mode matrix product of $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is denoted by $\mathcal{A}_{(n)}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n}\times G_{n+1}\times\ldots\times G_{N}}$ contains the element of $a_{g_{1},g_{2},\ldots,g_{n}}$ at $g_{n}$-th row and at $j$th column, denoted as
$j=1+\sum_{\ell=1,\ell\neq n}^{N}(g_{\ell}-1)\prod_{m=1,m\neq n}^{\ell-1}G_{m}$ (9)
Definition 4: The $n$-mode product between tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ and a matrix $L\in\mathbb{R}^{J_{n}\times G_{n}}$ is defined as $\mathcal{B}=\mathcal{A}\times_{n}L$, where the product of tensor $\mathcal{B}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n}\times G_{n+1}\times\ldots\times G_{N}}$. In terms of element wise, the $n$-mode product of tensor $\mathcal{B}$ can be computed as
$b_{g_{1}g_{2}\ldots g_{n-1}j_{n}g_{n+1}\ldots g_{N}}=\sum_{g_{1}=1}^{g_{n}}a_{g_{1}g_{2}\ldots g_{n}\ldots g_{N}}b_{j_{n}g_{n}}$ (10)
Definition 5: The tensor inner product between tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ and $\mathcal{B}\in\mathbb{R}^{J_{1}\times J_{2}\times\ldots\times J_{m}\times\ldots\times J_{M}}$ is denoted as
$\mathcal{C}=\textlangle\mathcal{A},\mathcal{B}\textrangle_{n,m}$ (11)
The element wise of tensor inner product $\mathcal{A}$ and $\mathcal{B}$ can be written explicitly as below
$c_{g_{1}g_{2}\ldots g_{n-1}g_{n+1}\ldots g_{N}j_{1}j_{2}\ldots j_{m-1}j_{m+1}\ldots j_{M}}=\sum_{q_{1}=1}^{q_{p}}a_{g_{1}g_{2}\ldots g_{n-1}g_{q}g_{n+1}\ldots g_{N}}b_{j_{1}j_{2}\ldots j_{m-1}j_{q}j_{m+1}\ldots j_{M}}$ (12)
where $i_{n}=j_{m}=q_{p}$ and the tensor $\mathcal{C}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n+1}\times\ldots\times G_{N}\times J_{1}\times J_{2}\times\ldots\times J_{m-1}\times J_{m+1}\times\ldots\times J_{M}}$.
Definition 6: The tensor scalar product of tensors $\mathcal{A},\mathcal{B}\in\coprod^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n+1}\times\ldots\times G_{N}}$ is the inner product of tensor $\mathcal{A}$ and $\mathcal{B}$, denoted as below
$\langle{\mathcal{A},\mathcal{B}}\rangle\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\sum_{g_{1}=1}^{G_{1}}\sum_{g_{2}=1}^{G_{2}}\ldots\sum_{g_{n}=1}^{G_{N}}a_{g_{1}g_{2}\ldots g_{n-1}g_{n+1}\ldots g_{N}}b_{g_{1}g_{2}\ldots g_{n-1}g_{n+1}\ldots g_{N}}$ (13)
Definition 7: The Frobenius norm of tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ also known as the “energy” of the tensor, defined as
$\|\mathcal{A}\|_{F}=\varepsilon_{\mathcal{A}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\sqrt{\langle{\mathcal{A},\mathcal{A}}\rangle}$ (14)
3.5.2 The Higher Order Singular Value Decomposition (HOSVD)
Tensor factorization is a method for reducing a large volume of data to a smaller set of data points. This is achieved by leveraging the collinearity of its heterogeneous data [wang2017tensor]. CANDECOMP/PARAFAC is a technique for decomposing data into its constituent parts. It is one of the most widely used tensor decomposition techniques, as referenced in the works of Kolda et al. (2009) [kolda2009tensor]. and Kiers (2000) [kiers2000towards]. Conversely, the HOSVD represents an extension of the matrix singular value decomposition (SVD). In order to provide a more thorough explanation of the higher-order SVD (HOSVD), this paper will draw upon certain concepts from Wang and Song’s work on two-mode tensor orders $N$ [wang2017tensor].
3.5.3 Tensor Distance
We know that, the Euclidean distance between data points $x_{i}$ and cluster centers $a_{k}$ can be represented as:
$d(x_{i},a_{k})=\|x_{i}-a_{k}\|^{2}$ (15)
Definition 8: The tensor distance of a given $\mathcal{X},\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{N}}$ $(N\geq 1)$, and $x$, $a$ denoted as the vector representation of tensor $\mathcal{X},\mathcal{A}$, such as the element wise of tensor $\mathcal{XA}_{G_{1}\times G_{2}\times\ldots\times G_{N}}$ $(1\leq g_{i}\leq G_{n},1\leq i\leq N)$ in $\mathcal{X},\mathcal{A}$ is corresponding to $x_{l}$, $a_{l}$, i.e., the $l$-th element in $x$, the $l$-th element in $a$, where $l=1+\sum_{i^{\prime}=1,i^{\prime}\neq i}^{G_{1}\times G_{2}\times\ldots\times G_{N}}(g_{l}-1)\prod_{m=1,m\neq i}^{i^{\prime}-1}G_{m}$. Then the tensor distance (TD) between $\mathcal{X}$ and $\mathcal{A}$ can be defined as
$d_{TD}(\mathcal{X},\mathcal{A})=\sqrt{\sum_{l,m=1}^{G_{1}\times G_{2}\times\ldots\times G_{N}}\rho_{lm}(x_{l}-a_{l})(x_{m}-a_{m})}=\sqrt{(x-a)^{T}P(x-a)}$ (16)
where $\rho_{lm}$ is the metric coefficient defined by
$\rho_{lm}=\frac{1}{2\pi\theta}\frac{\text{ exp }\big\{\|q_{1}-q_{m}\|_{2}^{2}\big\}}{2\theta}$ (17)
and $\|q_{1}-q_{m}\|_{2}$ is defined as the location distance between $\mathcal{X}_{g_{1}g_{2}\ldots g_{N}}$ and $\mathcal{A}_{g_{1}g_{2}\ldots g_{N}}$
$\|q_{1}-q_{m}\|_{2}=\sqrt{(g_{1}+g_{1}^{\prime})^{2}+(g_{2}+g_{2}^{\prime})^{2}+\ldots+(g_{n}+g_{n}^{\prime})^{2}+\ldots+(g_{N}+g_{N}^{\prime})^{2}}$ (18)
Here, $\theta$ is a user-defined parameter with $\theta>0$ and $P$ is the metric matrix.
3.6 Summary and Research Gaps
A review of extant literature reveals several significant research gaps that motivate the proposed approach, including the limited integration of research findings. While federated learning, multi-view clustering, and tensor decomposition have been the focus of extensive research study, their integration remains largely unexplored. Existing federated clustering methods do not leverage multi-view structure, and multi-view clustering approaches are not designed for federated environments. In the context of communication inefficiency, contemporary federated multi-view methods do not leverage tensor structure for the purpose of communication compression. The high-dimensional nature of multi-view data exacerbates communication overhead in federated settings, requiring novel tensor-aware compression schemes.
Existing privacy-preserving mechanisms are not optimized for multi-view tensor data, potentially leading to information leakage across views or reduced clustering utility. The development of novel differential privacy mechanisms tailored to tensor structures is imperative. A further issue is the absence of personalization. The prevailing assumption underlying most federated clustering approaches is that data distributions across clients are homogeneous. In practice, clients frequently possess heterogeneous data characteristics, necessitating customized clustering solutions that adapt to local data patterns while leveraging global knowledge. As indicated in the theoretical understanding, the convergence properties and privacy guarantees of federated multi-view tensor clustering remain to be fully elucidated. A theoretical analysis is necessary to provide convergence guarantees and characterize the privacy-utility trade-offs.
The proposed frameworks address these limitations by integrating advanced tensor decomposition with personalized federated learning, leveraging a heat-kernel-enhanced tensorized approach. The aforementioned framework provides both theoretical guarantees and practical advantages.
4 Proposed Methodology
This section presents the proposed heat-kernel enhanced tensorized framework for personalized federated multi-view clustering. The initial step involves formulating the problem setting and introducing the federated kernel Euclidean distance metric. In the following section, the calculation of federated heat-kernel coefficients adapted to the multi-view tensor data structure will be delineated. Finally, the overall algorithmic framework must be described, including local updates, global aggregation, and privacy-preserving mechanisms.
4.1 Problem Formulation
Let $\mathcal{L}=\{L_{1},L_{2},\ldots,L_{M}\}$ be a set of $M$ clients in the federated system. Each client $L_{\ell}$ possesses local multi-view data that can be organized into tensor form $\mathcal{X}_{[\ell]}\in\mathbb{R}^{n(\ell)\times d_{1}^{(\ell)}\times d_{2}^{(\ell)}\times\cdots\times d_{s(\ell)}^{(\ell)}}$, where:
  • $n(\ell)$ is the number of samples at client $\ell$
  • $s(\ell)$ is the number of views at client $\ell$
  • $d_{h}^{(\ell)}$ is the feature dimension of view $h$ at client $\ell$
Alternatively, we can represent the data as a collection of view-specific matrices $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ where $X_{[\ell]}^{h}\in\mathbb{R}^{n(\ell)\times d_{h}^{(\ell)}}$, which can be tensorized for advanced decomposition techniques.
4.2 Federated Kernel Euclidean Distance
In federated environment, we assume participated clients shared different number of data but the same feature component. Thus, the ground truth or predicted clusters of multi-view data sets stored on their data centers may vary. In this way, each data points on clients will shape a unified global multi-view datasets in a certain problem or case. Therefore, it is explained a lot the definition of “heterogeneous” on data. This characteristic is denoted by $c(\ell)$, representing the number of clusters of multi-view data sets held by the $\ell$-th client. By following this direction, we extend the heat-kernel enhanced distance metric KED in Eq. 4 to operate in federated environment. For client $\ell$, the kernel Euclidean distance (KED) between $x_{[\ell]ij}^{h}$ (representing data point $i$, $\forall i=1,\ldots,n(\ell)$) and cluster center tensor $a_{[\ell]kj}^{h}$ (representing cluster $k$, $\forall k=1,\ldots,c(\ell)$) is:
$\text{FKED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})=1-\exp\left(-\sum_{h=1}^{s(\ell)}\sum_{j=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij}^{h}\left(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h}\right)^{2}\right)$ (19)
where $\delta_{[\ell]ij}^{h}$ are heat-kernel coefficients held by $M$ clients.
4.3 Federated Heat-Kernel Coefficients
We focus on developing federated heat-kernel coefficients (FedH-KC), which are computed locally by each client to transform feature distances into privacy-preserving exponential kernel measures tailored to the federated setting. For the $j$-th feature in view $h$, we define FedH-KC $\delta_{\left[\ell\right]ij}^{h}$ using two distinct estimators.
$\text{FedH-KC}=\delta_{\left[\ell\right]ij}^{h}=\frac{x_{\left[\ell\right]ij}^{h}-\min_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)}{\max_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)-\min_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)+\epsilon}\qquad\forall j,h,\ell$ (20)
where $\epsilon>0$ prevents division by zero when features have identical values. This normalizes features within each view for consistent distance calculations across clients with different data distributions. and
$\text{FedH-KC}=\delta_{\left[\ell\right]ij}^{h}=\left|{x_{\left[\ell\right]ij}^{h}-{{\bar{x}}_{\left[\ell\right]j}}^{h}}\right|\qquad\forall j,h,\ell$ (21)
where $\bar{x}_{[\ell]j}^{h}=\frac{1}{n(\ell)}\sum_{i=1}^{n(\ell)}x_{[\ell]ij}^{h}$ represents the mean value of the $j$-th feature in view $h$ for client $\ell$, harmonizing with the centralized case notation.
5 The Enhanced Federated Kernel Multi-View Clustering
In this section, the proposed Enhanced Federated Kernel Multi-View Clustering (E-FKMVC) algorithm will be presented. The E-FKMVC algorithm is a refinement of the traditional multi-view clustering framework that operates within a federated learning environment. This refinement enables multiple clients to collaboratively cluster their local multi-view datasets while preserving data privacy.
5.1 Personalized Federated Learning
Considers there are $M$ number of clients storing multi-view data sets $X_{[\ell]}=X_{[\ell]i}^{h}$ for all $\ell=1,\ldots,M$ and $h=1,\ldots,s(\ell)$. Here, we assume each client is storing a different number of dimensionality of data points $x_{i}$. To solve clustering problem for these multi-view data sets distributed on $M$ clients, we propose an efficient federated kernel multi-view clustering (E-FKMVC). The proposed E-FKMVC modified Eq. 6 in the following way
$J_{E-FKMVC}^{(\ell)}=J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})=\sum_{\ell=1}^{M}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{~FKED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})$ (22)
subject to:
$\sum_{k=1}^{c}\mu_{[\ell]ik}^{*}=1,\mu_{[\ell]ik}^{*}\in[0,1]$ (23)
$\sum_{h=1}^{s}v_{[\ell]h}=1,v_{[\ell]h}\in[0,1]$ (24)
Due to the non-convex nature of the E-FKMVC objective function in Eq. 22, direct minimization is not feasible. We employ an alternating optimization strategy that iteratively updates the membership matrix $U_{[\ell]}^{*}$, view weights $V_{[\ell]}$, and cluster centers $A_{[\ell]}^{h}$ until convergence.
To derive the update rules, we formulate the Lagrangian incorporating the normalization constraints:
$\begin{split}\tilde{J}_{E-FKMVC}^{(\ell)}&=J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})\\ &\quad+\sum_{i=1}^{n(\ell)}{\lambda_{1i}}\left(\sum_{k=1}^{c(\ell)}{\mu_{[\ell]ik}^{*}-1}\right)+{\lambda_{2}}\left(\sum_{h=1}^{s(\ell)}{v_{[\ell]h}-1}\right)\end{split}$ (25)
where $\lambda_{1i}$} and $\lambda_{2}$} are Lagrange multipliers enforcing the fuzzy clustering constraints: membership normalization and view weight normalization, respectively.
Theorem 1 (E-FKMVC Update Rules).
For the federated objective function $J_{E-FKMVC}$ defined in Eq. 22, the necessary conditions for optimality yield the following update rules for client $\ell$:
Membership Matrix Update:
$\mu_{\left[\ell\right]ik}^{*}=\frac{{{{\left({\sum\limits_{h=1}^{s\left(\ell\right)}{v_{\left[\ell\right]h}^{\alpha}{\rm{FKED}}\left({x_{\left[\ell\right]ij}^{h},a_{\left[\ell\right]kj}^{h}}\right)}}\right)}^{-{{\left({m-1}\right)}^{-1}}}}}}{{\sum\limits_{k^{\prime}=1}^{c\left(\ell\right)}{{{\left({\sum\limits_{h=1}^{s\left(\ell\right)}{v_{\left[\ell\right]h}^{\alpha}{\rm{FKED}}\left({x_{\left[\ell\right]ij}^{h},a_{\left[\ell\right]k^{\prime}j}^{h}}\right)}}\right)}^{-{{\left({m-1}\right)}^{-1}}}}}}}$ (26)
Cluster Centers Update:
$a_{\left[\ell\right]kj}^{h}=\frac{{\sum\limits_{i=1}^{n\left(\ell\right)}{{{\left({\mu_{\left[\ell\right]ik}^{*}}\right)}^{m}}v_{\left[\ell\right]h}^{\alpha}\exp\left({-\delta_{\left[\ell\right]ij}^{h}{{\left\|{x_{\left[\ell\right]i}^{h}-a_{\left[\ell\right]k}^{h}}\right\|}^{2}}}\right)}}}{{\sum\limits_{i=1}^{n\left(\ell\right)}{{{\left({\mu_{\left[\ell\right]ik}^{*}}\right)}^{m}}v_{\left[\ell\right]h}^{\alpha}\exp\left({-\delta_{\left[\ell\right]ij}^{h}{{\left\|{x_{\left[\ell\right]i}^{h}-a_{\left[\ell\right]k}^{h}}\right\|}^{2}}}\right)}}}x_{\left[\ell\right]ij}^{h}$ (27)
View Weights Update:
${v_{\left[\ell\right]h}}=\frac{{{{\left({\sum\limits_{i=1}^{n\left(\ell\right)}{\sum\limits_{k=1}^{c\left(\ell\right)}{{{\left({\mu_{\left[\ell\right]ik}^{*}}\right)}^{m}}{\rm{FKED}}\left({x_{\left[\ell\right]ij}^{h},a_{\left[\ell\right]kj}^{h}}\right)}}}\right)}^{-{{\left({\alpha-1}\right)}^{-1}}}}}}{{\sum\limits_{h^{\prime}=1}^{s\left(\ell\right)}{{{\left({\sum\limits_{i=1}^{n\left(\ell\right)}{\sum\limits_{k=1}^{c\left(\ell\right)}{{{\left({\mu_{\left[\ell\right]ik}^{*}}\right)}^{m}}{\rm{FKED}}\left({x_{\left[\ell\right]ij}^{h^{\prime}},a_{\left[\ell\right]kj}^{h^{\prime}}}\right)}}}\right)}^{-{{\left({\alpha-1}\right)}^{-1}}}}}}}$ (28)
where the updates are performed iteratively until convergence, subject to the normalization constraints in Eqs. 23 and 24.
Proof.
We establish the necessary optimality conditions for the E-FKMVC objective function using Lagrangian optimization and derive the closed-form update rules for each parameter set. The proof proceeds by applying the method of Lagrange multipliers to the constrained optimization problem and solving the resulting system of equations.
Part I: Membership Matrix Update Rule Derivation
To derive the update equation for membership coefficients $\mu_{[\ell]ik}^{*}$ for client $\ell$, we begin by computing the partial derivative of the Lagrangian in Eq. 25 with respect to $\mu_{[\ell]ik}^{*}$.
Taking the partial derivative with respect to $\mu_{[\ell]ik}^{*}$:
$\frac{\partial\tilde{J}_{E-FKMVC}^{(\ell)}}{\partial\mu_{[\ell]ik}^{*}}$ $=\frac{\partial}{\partial\mu_{[\ell]ik}^{*}}\left[\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right]+\lambda_{1i}$ (29)
$=m(\mu_{[\ell]ik}^{*})^{m-1}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})+\lambda_{1i}$ (30)
Setting this derivative equal to zero for optimality:
$m(\mu_{[\ell]ik}^{*})^{m-1}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})+\lambda_{1i}=0$ (31)
Solving for $\mu_{[\ell]ik}^{*}$:
$(\mu_{[\ell]ik}^{*})^{m-1}$ $=-\frac{\lambda_{1i}}{m\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})}$ (32)
$\mu_{[\ell]ik}^{*}$ $=\left(-\frac{\lambda_{1i}}{m}\right)^{1/(m-1)}\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(m-1)}$ (33)
To determine the Lagrange multiplier $\lambda_{1i}$, we apply the normalization constraint $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$:
$\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}$ $=\left(-\frac{\lambda_{1i}}{m}\right)^{1/(m-1)}\sum_{k=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(m-1)}=1$ (34)
Solving for the Lagrange multiplier:
$\left(-\frac{\lambda_{1i}}{m}\right)^{1/(m-1)}$ $=\frac{1}{\sum_{k^{\prime}=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k^{\prime}}^{h})\right)^{-1/(m-1)}}$ (35)
Substituting this back into Eq. 33:
$\mu_{[\ell]ik}^{*}=\frac{\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(m-1)}}{\sum_{k^{\prime}=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k^{\prime}}^{h})\right)^{-1/(m-1)}}$ (36)
This establishes the membership matrix update rule in Eq. 26.
Part II: Cluster Centers Update Rule Derivation
For cluster center updates $a_{[\ell]kj}^{h}$, we differentiate the objective function in Eq. 22 with respect to $a_{[\ell]kj}^{h}$:
$\frac{\partial J_{E-FKMVC}^{(\ell)}}{\partial a_{[\ell]kj}^{h}}$ $=\frac{\partial}{\partial a_{[\ell]kj}^{h}}\left[v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right]$ (37)
$=v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\frac{\partial\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})}{\partial a_{[\ell]kj}^{h}}$ (38)
To compute $\frac{\partial\text{FKED}}{\partial a_{[\ell]kj}^{h}}$, we recall that:
$\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})=1-\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$ (39)
Let $\phi_{[\ell]ik}^{h}=\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}$. Then:
$\frac{\partial\text{FKED}}{\partial a_{[\ell]kj}^{h}}$ $=\frac{\partial}{\partial a_{[\ell]kj}^{h}}\left[1-\exp(-\phi_{[\ell]ik}^{h})\right]$ (40)
$=\exp(-\phi_{[\ell]ik}^{h})\frac{\partial\phi_{[\ell]ik}^{h}}{\partial a_{[\ell]kj}^{h}}$ (41)
$=\exp(-\phi_{[\ell]ik}^{h})\cdot 2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\cdot(-1)$ (42)
$=-2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$ (43)
Substituting this into Eq. 38 and setting equal to zero for optimality:
$v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\left[-2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)\right]=0$ (44)
Simplifying and rearranging:
$\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\delta_{[\ell]ij}^{h}$ $\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})=0$ (45)
Let $w_{[\ell]ijk}^{h}=(\mu_{[\ell]ik}^{*})^{m}v_{[\ell]h}^{\alpha}\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$. Then:
$\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})$ $=0$ (46)
$\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}x_{[\ell]ij}^{h}$ $=a_{[\ell]kj}^{h}\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}$ (47)
$a_{[\ell]kj}^{h}$ $=\frac{\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}x_{[\ell]ij}^{h}}{\sum_{i=1}^{n(\ell)}w_{[\ell]ijk}^{h}}$ (48)
This yields the cluster center update rule in Eq. 27.
Part III: View Weights Update Rule Derivation
For view weight updates $v_{[\ell]h}$, we differentiate the Lagrangian with respect to $v_{[\ell]h}$:
$\frac{\partial\tilde{J}_{E-FKMVC}^{(\ell)}}{\partial v_{[\ell]h}}$ $=\frac{\partial}{\partial v_{[\ell]h}}\left[\sum_{h^{\prime}=1}^{s(\ell)}v_{[\ell]h^{\prime}}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h^{\prime}},a_{[\ell]k}^{h^{\prime}})\right]+\lambda_{2}$ (49)
$=\alpha v_{[\ell]h}^{\alpha-1}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})+\lambda_{2}$ (50)
Setting this derivative equal to zero for optimality:
$\alpha v_{[\ell]h}^{\alpha-1}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})+\lambda_{2}=0$ (51)
Solving for $v_{[\ell]h}$:
$v_{[\ell]h}^{\alpha-1}$ $=-\frac{\lambda_{2}}{\alpha\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})}$ (52)
$v_{[\ell]h}$ $=\left(-\frac{\lambda_{2}}{\alpha}\right)^{1/(\alpha-1)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(\alpha-1)}$ (53)
Applying the normalization constraint $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$:
$\sum_{h=1}^{s(\ell)}v_{[\ell]h}$ $=\left(-\frac{\lambda_{2}}{\alpha}\right)^{1/(\alpha-1)}\sum_{h=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(\alpha-1)}=1$ (54)
Solving for the Lagrange multiplier:
$\left(-\frac{\lambda_{2}}{\alpha}\right)^{1/(\alpha-1)}$ $=\frac{1}{\sum_{h^{\prime}=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h^{\prime}},a_{[\ell]k}^{h^{\prime}})\right)^{-1/(\alpha-1)}}$ (55)
Substituting back:
$v_{[\ell]h}=\frac{\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right)^{-1/(\alpha-1)}}{\sum_{h^{\prime}=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h^{\prime}},a_{[\ell]k}^{h^{\prime}})\right)^{-1/(\alpha-1)}}$ (56)
This establishes the view weight update rule in Eq. 28.
Part IV: Convergence Analysis and Optimality Conditions
The derived update rules satisfy the Karush-Kuhn-Tucker (KKT) conditions for the constrained optimization problem. Specifically:
  1. 1.
    Stationarity: $\nabla J_{E-FKMVC}^{(\ell)}+\sum_{i=1}^{n(\ell)}\lambda_{1i}\nabla g_{i}+\lambda_{2}\nabla h=0$, where $g_{i}(\mu)=\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}-1$ and $h(v)=\sum_{h=1}^{s(\ell)}v_{[\ell]h}-1$.
  2. 2.
    Primal feasibility: $g_{i}(\mu)=0$ and $h(v)=0$ for all $i$.
  3. 3.
    Dual feasibility: All Lagrange multipliers are finite and well-defined.
The alternating optimization scheme converges to a local optimum under the following conditions:
  • Boundedness and Continuity: The objective function $J_{E-FKMVC}^{(\ell)}$ is continuous and bounded below. This follows from the fact that FKED values are bounded in $[0,1]$ by construction (Eq. 4), ensuring the weighted sum in the objective function remains finite. The exponential decay in the heat kernel guarantees that distances approach finite limits as data points move infinitely far apart.
  • Constraint Compactness: The constraint set is compact and convex. The membership matrix constraints $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$ with $\mu_{[\ell]ik}^{*}\in[0,1]$ define probability simplices, which are compact convex sets in $\mathbb{R}^{c(\ell)}$. Similarly, the view weight constraints $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$ with $v_{[\ell]h}\in[0,1]$ form a compact convex simplex in $\mathbb{R}^{s(\ell)}$. The cluster center parameters are unconstrained but remain bounded due to the data distribution.
  • Unique Subproblem Solutions: Each subproblem (updating $U$, AA, or $V$ while fixing others) has a unique solution. For membership updates, the strictly positive denominators in Eq. 26 ensure uniqueness since FKED values are always positive for distinct data points and cluster centers. For cluster center updates, the weighted exponential terms in Eq. 27 create a strictly convex weighted least-squares problem with a unique minimum. For view weight updates, the strictly decreasing nature of the power function with exponent $1/(\alpha-1)$ in Eq. 28 guarantees uniqueness of the normalized solution.
These conditions are satisfied by construction, as the FKED function is continuous and bounded, the probability simplex constraints define compact convex sets, and the update rules yield unique solutions when the denominators are non-zero (guaranteed by the strict positivity of FKED values and the non-degeneracy assumption that no data points are identical to cluster centers).
Algorithm 1 summarizes the steps of the proposed E-FKMVC algorithm. This algorithm iteratively updates the membership matrices, cluster centers, and view weights for each client until convergence is achieved.
Algorithm 1 Enhanced Federated Kernel Multi-View Clustering (E-FKMVC)
0: Multi-view data $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ for clients $\ell=1,\ldots,M$, number of clusters $c(\ell)$, fuzziness parameter $m>1$, view weight exponent $\alpha>1$, heat-kernel parameters $\delta_{[\ell]ij}^{h}$
0: Membership matrices $U_{[\ell]}^{*}$, cluster centers $A_{[\ell]}^{h}$, view weights $V_{[\ell]}$
1:Initialization: Randomly initialize $U_{[\ell]}^{*}$, $A_{[\ell]}^{h}$, and $V_{[\ell]}$ for all clients $\ell$
2:repeat
3:  for each client $\ell=1$ to $M$ do
4:   Update membership matrix $U_{[\ell]}^{*}$ using Eq. 26
5:   Update cluster centers $A_{[\ell]}^{h}$ using Eq. 27
6:   Update view weights $V_{[\ell]}$ using Eq. 28
7:  end for
8:until convergence criteria met (e.g., changes in $U_{[\ell]}^{*}$, $A_{[\ell]}^{h}$, and $V_{[\ell]}$ below a threshold)
6 Proposed Federated Tensorized Clustering
In this section, an introduction is provided to the proposed federated tensorized clustering framework, which integrates heat-kernel-enhanced multi-view clustering with tensor decomposition techniques. This integration enables the efficient management of high-dimensional multi-view data distributed across multiple clients. The central concept is to utilize tensor representations to capture the multi-modal structure of the data while leveraging heat-kernel based distance metrics to enhance clustering performance in a federated learning setting.
6.1 Tensorized Heat-Kernel Enhanced Local Clustering
Each client $\ell$ possesses multi-view data $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$, where $X_{[\ell]}^{h}\in\mathbb{R}^{n(\ell)\times d_{h}^{(\ell)}}$ represents the data matrix for view $h$ with $n(\ell)$ samples and $d_{h}^{(\ell)}$ features. To effectively capture the multi-view structure, we propose a tensorized heat-kernel enhanced clustering approach. Two main components are involved: tensor representation of multi-view data and tensorized heat-kernel distance metric. Decomposition methods are then applied to represent cluster centers efficiently.
6.1.1 Tensor Data Organization
As multi-view data inherently possesses a multi-modal structure, we organize the data into a tensor format. The multi-view data from client $\ell$ is represented as a collection of matrices $\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$. To facilitate tensor operations, we stack these matrices along a new mode, resulting in a 3-mode tensor representation. The 3-mode tensor captures the relationships across samples, features, and views simultaneously. Mathematically, the multi-view data at client $\ell$ is organized into a 3-mode tensor as follows:
$\mathcal{X}_{[\ell]}\in\mathbb{R}^{n(\ell)\times D_{[\ell]}\times s(\ell)}$ (57)
where each slice $\mathcal{X}_{[\ell],:,:,h}=X_{[\ell]}^{h}$ corresponds to the data matrix of view $h$, and $D_{[\ell]}$ is the maximum feature dimension across all views for client $\ell$ ($D_{[\ell]}=\max_{h}d_{h}^{(\ell)}$). If views have varying feature dimensions, zero-padding is applied to ensure consistent tensor dimensions. Specifically, for view $h$ with $d_{h}^{(\ell)}$ features, we define:
$\mathcal{X}_{[\ell],i,j,h}=\begin{cases}X_{[\ell],i,j}^{h}&\text{if }j\leq d_{h}^{(\ell)}\\ 0&\text{if }j>d_{h}^{(\ell)}\end{cases}$ (58)
Thus, the complete tensor representation is given by:
$\mathcal{X}_{[\ell]}=\left[\mathcal{X}_{[\ell],:,:,1}|\mathcal{X}_{[\ell],:,:,2}|\ldots|\mathcal{X}_{[\ell],:,:,s(\ell)}\right]\in\mathbb{R}^{n(\ell)\times D_{[\ell]}\times s(\ell)}$ (59)
This tensorization enables a unified representation of multi-view data, allowing for efficient processing and analysis of the inherent multi-modal relationships. The tensor structure facilitates the application of tensor decomposition techniques for dimensionality reduction and feature extraction, which are crucial for effective clustering in high-dimensional spaces. Key advantages of this tensor representation include:
  • Preservation of multi-view relationships
  • Facilitation of tensor-based operations
  • Enabling advanced decomposition methods for clustering
6.1.2 Tensorized Kernel Euclidean Distance
To enable effective clustering in the tensorized multi-view space, we extend the heat-kernel enhanced distance metric to operate on tensor slices. The traditional heat-kernel enhanced distance metric (KED) is defined for vector data points. In our tensorized setting, we adapt this metric to measure distances between tensor slices corresponding to different views. The idea is to compute the distance between tensor slices while incorporating heat-kernel weighting to emphasize local structures in the data.
The tensorized kernel Euclidean distance (TKED) captures the similarity between tensor slices while incorporating the heat-kernel weighting to emphasize local structures in the data. Based on the original KED definition in Eq. 4, the TKED is formulated as follows:
$\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h},\mathcal{A}_{[\ell],k,:}^{h})=1-\exp\left(-\sum_{h=1}^{s(\ell)}\sum_{j=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij}^{h}\left(\mathcal{X}_{[\ell],i,j}^{h}-\mathcal{A}_{[\ell],k,j}^{h}\right)^{2}\right)$ (60)
where $\mathcal{A}_{[\ell]}\in\mathbb{R}^{c(\ell)\times D_{[\ell]}\times s(\ell)}$ represents the cluster center tensor for client $\ell$, with each slice $\mathcal{A}_{[\ell],k,:}^{h}$ corresponding to the cluster center for cluster $k$ in view $h$. The heat-kernel coefficients $\delta_{[\ell]ij}^{h}$ are computed similarly to Eq. 20, but adapted for tensor slices. Specifically, the heat-kernel coefficients for the tensorized setting are defined as:
$\delta_{[\ell]ij}^{h}=\frac{\mathcal{X}_{[\ell],i,j}^{h}-\min_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}}{\max_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}-\min_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}}$ (61)
As in the vector case, the heat-kernel coefficients $\delta_{[\ell]ij}^{h}$ are computed based on the normalized feature values across samples for each feature dimension $j$ in view $h$. This normalization ensures that the heat-kernel weighting effectively captures local structures in the tensorized multi-view space. The TKED metric thus provides a robust measure of similarity between tensor slices, facilitating effective clustering in the federated tensorized multi-view setting.
6.1.3 Tensor Decomposition for Cluster Centers
To reduce computational complexity and enable efficient communication, we represent cluster centers using a tensor decomposition approach. Specifically, we employ the PARAFAC2 decomposition to model the cluster center tensor $\mathcal{A}_{[\ell]}$ as:
$\mathcal{A}_{[\ell]}=\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$ (62)
where $\mathcal{G}_{[\ell]}$ is the core tensor and $\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ are factor matrices corresponding to modes 1, 2, and 3, respectively.
PARAFAC2 is chosen for its flexibility in handling varying dimensions across one mode (samples) while maintaining consistent factor matrices in the other modes (features and views). This property is particularly advantageous in federated settings where clients may have heterogeneous data distributions. However, other tensor decomposition methods (e.g., Tucker, CP) can also be adapted based on specific application requirements.
The Tucker decomposition could be employed when the data exhibits strong multilinear interactions across all modes, while CP decomposition may be preferred for its simplicity and interpretability in scenarios with more uniform data structures. The Tucker decomposition allows for a more compact representation by capturing interactions through a core tensor, which can be beneficial when the data has complex relationships. CP decomposition, on the other hand, provides a straightforward sum of rank-one tensors, making it easier to implement and interpret in certain contexts. In this work, we also focus on Tucker decomposition as an alternative to PARAFAC2, providing flexibility in modeling diverse data characteristics across clients.
The representation cluster centers using Tucker decomposition is given by:
$\mathcal{A}_{[\ell]}=\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$ (63)
where $\mathcal{G}_{[\ell]}$ is the core tensor and $\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$ are factor matrices corresponding to modes 1, 2, and 3, respectively.
6.2 The Proposed FedHK-PARAFAC2 Local Objective Function
The FedHK-PARAFAC2 local objective function for client $\ell$ is defined as:
$J_{FedHK-PARAFAC2}^{(\ell)}=\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})$ (64)
subject to the constraints:
$\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}$ $=1,\quad\mu_{[\ell]ik}^{*}\in[0,1]$ (65)
$\sum_{h=1}^{s(\ell)}v_{[\ell]h}$ $=1,\quad v_{[\ell]h}\in[0,1]$ (66)
$\mathcal{A}_{[\ell]}$ $=\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$ (67)
Theorem 2 (The FedHK-PARAFAC2 Update Rules).
The optimal updates for the tensorized federated heat-kernel enhanced multi-view clustering at client $\ell$ are given by:
Membership updates:
$\mu_{[\ell]ik}^{*}=\frac{\left(\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})\right)^{-\frac{1}{m-1}}}{\sum_{k^{\prime}=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k^{\prime},:}^{h})\right)^{-\frac{1}{m-1}}}$ (68)
Core tensor updates:
$\mathcal{G}_{[\ell]}=\arg\min_{\mathcal{G}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}\times_{3}\mathbf{R})_{[\ell],k,:}^{h}\|^{2}$ (69)
View weight updates:
$v_{[\ell]h}=\frac{\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})\right)^{-\frac{1}{\alpha-1}}}{\sum_{h^{\prime}=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h^{\prime}}-\mathcal{A}_{[\ell],k,:}^{h^{\prime}})\right)^{-\frac{1}{\alpha-1}}}$ (70)
Cluster center updates via PARAFAC2 decomposition:
$\mathcal{A}_{[\ell]}=\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$ (71)
Factor matrix updates:
$\mathbf{P}_{[\ell]}$ $=\arg\min_{\mathbf{P}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}\times_{3}\mathbf{R})_{[\ell],k,:}^{h}\|^{2}$ (72)
$\mathbf{Q}_{[\ell]}$ $=\arg\min_{\mathbf{Q}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}\times_{3}\mathbf{R})_{[\ell],k,:}^{h}\|^{2}$ (73)
$\mathbf{R}_{[\ell]}$ $=\arg\min_{\mathbf{R}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}\times_{3}\mathbf{R})_{[\ell],k,:}^{h}\|^{2}$ (74)
The Proof of Theorem 2.
By applying the method of Lagrange multipliers to the tensorized local objective in Eq. LABEL:eqn:tensor_local_obj and following similar derivation steps as in Theorem 1, we derive the update rules for membership, core tensor, factor matrices, and view weights. The key steps involve differentiating the objective with respect to each variable while enforcing the constraints, leading to the closed-form solutions presented in Eqs. 68 to 70. The alternating optimization approach ensures convergence to a local optimum under the same conditions outlined in Theorem 1.
First, we derive the membership update by differentiating the Lagrangian with respect to $\mu_{jk}^{(i)}$ and setting it to zero, leading to Eq. 68. Next, we derive the core tensor and factor matrix updates by fixing other variables and solving the resulting least-squares problems, yielding Eqs. 69 to 74. Finally, we derive the view weight update by differentiating with respect to $v_{h}^{(i)}$ and applying the normalization constraint, resulting in Eq. 70. Finally, we verify that these updates satisfy the KKT conditions and ensure convergence under the specified conditions.
6.3 The Proposed FedHK-Tucker Local Objective Function
By implementing Tucker decomposition in Eq. 63, the FedHK-Tucker local objective function for client $\ell$ is defined as:
$J_{FedHK-Tucker}^{(\ell)}=\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})$ (75)
subject to the constraints:
$\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}$ $=1,\quad\mu_{[\ell]ik}^{*}\in[0,1]$ (76)
$\sum_{h=1}^{s(\ell)}v_{[\ell]h}$ $=1,\quad v_{[\ell]h}\in[0,1]$ (77)
$\mathcal{A}_{[\ell]}$ $=\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$ (78)
Theorem 3 (The FedHK-Tucker Update Rules).
The optimal updates for the tensorized federated heat-kernel enhanced multi-view clustering with Tucker decomposition at client $\ell$ are given by: Membership updates:
$\mu_{[\ell]ik}^{*}=\frac{\left(\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})\right)^{-\frac{1}{m-1}}}{\sum_{k^{\prime}=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k^{\prime},:}^{h})\right)^{-\frac{1}{m-1}}}$ (79)
Core tensor updates:
$\mathcal{G}_{[\ell]}=\arg\min_{\mathcal{G}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{U}\times_{2}\mathbf{V}\times_{3}\mathbf{W})_{[\ell],k,:}^{h}\|^{2}$ (80)
View weight updates:
$v_{[\ell]h}=\frac{\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})\right)^{-\frac{1}{\alpha-1}}}{\sum_{h^{\prime}=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h^{\prime}}-\mathcal{A}_{[\ell],k,:}^{h^{\prime}})\right)^{-\frac{1}{\alpha-1}}}$ (81)
Cluster center updates via Tucker decomposition:
$\mathcal{A}_{[\ell]}=\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$ (82)
Factor matrix updates:
$\mathbf{U}_{[\ell]}$ $=\arg\min_{\mathbf{U}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{U}\times_{2}\mathbf{V}\times_{3}\mathbf{W})_{[\ell],k,:}^{h}\|^{2}$ (83)
$\mathbf{V}_{[\ell]}$ $=\arg\min_{\mathbf{V}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{U}\times_{2}\mathbf{V}\times_{3}\mathbf{W})_{[\ell],k,:}^{h}\|^{2}$ (84)
$\mathbf{W}_{[\ell]}$ $=\arg\min_{\mathbf{W}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{U}\times_{2}\mathbf{V}\times_{3}\mathbf{W})_{[\ell],k,:}^{h}\|^{2}$ (85)
The Proof of Theorem 3.
The proof follows similar steps as in Theorem 2, with adjustments for the Tucker decomposition structure. By applying the method of Lagrange multipliers to the Tucker-based local objective in Eq. 75 and differentiating with respect to each variable while enforcing the constraints, we derive the closed-form solutions presented in Eqs. 79 to 81. The alternating optimization approach ensures convergence to a local optimum under the same conditions outlined in Theorem 1. ∎
6.4 Local Clustering Algorithms
Provided below the differences between the local clustering algorithms for E-FKMVC and the tensorized FedHK-PARAFAC2/FedHK-Tucker methods. Table 1 summarizes the key distinctions.
Table 1: Comparison of Local Clustering Algorithms
Aspect E-FKMVC FedHK-PARAFAC2/FedHK-Tucker
Data Structure Multi-view matrices Multi-view tensors
Distance Metric FKED (Eq. 4) TKED (Eq. 60)
Cluster Center Representation Direct centers $A^{h}$ Decomposed tensors $\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ or $\mathcal{G}_{[\ell]},\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$
Update Rules Eqs. 26 - 28 Eqs. 68 - 70 or 79 - 81
Optimization Method Direct updates Alternating Least Squares (ALS) for tensor factors
As illustrated in Table 1, the E-FKMVC algorithm operates directly on multi-view data matrices using the FKED metric, while the tensorized methods utilize multi-view tensors and the TKED metric. The cluster centers in E-FKMVC are represented directly, whereas in the tensorized methods, they are represented via tensor decompositions (PARAFAC2 or Tucker). The update rules differ accordingly, with the tensorized methods employing Alternating Least Squares (ALS) for updating tensor factors.
The PARAFAC2 and Tucker decompositions in E-FKMVC enable efficient representation and updating of cluster centers in high-dimensional multi-view spaces, facilitating effective clustering in federated settings. For both tensorized methods, the local clustering algorithms involve iterative updates of memberships, tensor factors, and view weights until convergence. Table 2 summarizes the key differences between the FedHK-PARAFAC2 and FedHK-Tucker methods.
Table 2: Comparison of FedHK-PARAFAC2 and FedHK-Tucker Methods
Aspect FedHK-PARAFAC2 FedHK-Tucker
Decomposition Type PARAFAC2 Tucker
Cluster Center Representation $\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ $\mathcal{G}_{[\ell]},\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$
Update Rules Eqs. 68 - 70 Eqs. 79 - 81
Optimization Method ALS for PARAFAC2 factors ALS for Tucker factors
As can be seen in Table 2, the primary difference between the two methods lies in the type of tensor decomposition employed (PARAFAC2 vs. Tucker) and the corresponding representation of cluster centers. The update rules and optimization methods are adapted accordingly to suit the specific decomposition framework used in each method. See Algorithm 2 and Algorithm 3 for the detailed local clustering procedures for FedHK-PARAFAC2 and FedHK-Tucker, respectively.
ALS Optimization for PARAFAC2 Factors
The Alternating Least Squares (ALS) optimization for updating the PARAFAC2 factors involves iteratively optimizing each factor matrix while keeping the others fixed. The update steps for the factor matrices $\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ are derived by solving least-squares problems based on the current estimates of the other factors and the core tensor $\mathcal{G}_{[\ell]}$. This process continues until convergence, ensuring that the cluster center tensor $\mathcal{A}_{[\ell]}$ is accurately represented in the decomposed form.
ALS Optimization for Tucker Factors
Similar to the PARAFAC2 case, the ALS optimization for Tucker factors involves iteratively updating each factor matrix $\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$ while keeping the others fixed. The updates are obtained by solving least-squares problems that minimize the reconstruction error of the cluster center tensor $\mathcal{A}_{[\ell]}$ based on the current estimates of the core tensor $\mathcal{G}_{[\ell]}$ and the other factor matrices. This iterative process continues until convergence, ensuring an accurate representation of the cluster centers in the Tucker decomposed form.
For both tensorized methods, the ALS optimization provides an efficient means of updating the tensor factors, enabling effective clustering in high-dimensional multi-view spaces while leveraging the benefits of tensor decomposition. Please refer to previous sections for the specific update equations and procedures involved in the ALS optimization for both PARAFAC2 and Tucker decompositions.
Algorithm 2 Tensorized Federated Heat-Kernel Enhanced Multi-View Clustering (Local FedHK-PARAFAC2)
Input: Local tensor data $\mathcal{X}_{[\ell]}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{P}_{g},\mathbf{Q}_{g},\mathbf{R}_{g}\}$, personalization parameters $\{\lambda_{[\ell]},\rho_{[\ell]}\}$, $g$ is the global model index
Initialize:
$\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{[\ell]}\mathcal{G}_{g}+(1-\lambda_{[\ell]})\mathcal{G}_{[\ell]}$
$\mathbf{P}_{[\ell]}^{(0)},\mathbf{Q}_{[\ell]}^{(0)},\mathbf{R}_{[\ell]}^{(0)}\leftarrow$ similarly personalized initialization
$v_{[\ell]}^{(0)}\leftarrow\frac{1}{s(\ell)}\mathbf{1}$
$t\leftarrow 0$
repeat
  $t\leftarrow t+1$
  // Compute tensorized heat-kernel coefficients
  Calculate $\delta_{[\ell]ij}^{h}$ using Eq. 61
  // Reconstruct cluster centers from decomposition
  $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$
  // Update memberships
  Update $U_{[\ell]}^{*^{(t)}}$ using Eq. 68
  // Update tensor factors using ALS
  Update $\mathcal{G}_{[\ell]}^{(t)}$ using Eq. 69
  Update $\mathbf{P}_{[\ell]}^{(t)}$ using Eq. 72
  Update $\mathbf{Q}_{[\ell]}^{(t)}$ using Eq. 73
  Update $\mathbf{R}_{[\ell]}^{(t)}$ using Eq. 74
  // Update view weights
  Update $v_{[\ell]}^{(t)}$ using Eq. 70
  // Check convergence
  Compute $J_{[\ell]}^{tensor}$ using Eq. 64
until convergence or max iterations
return $U_{[\ell]}^{*^{(t)}},\mathcal{G}_{[\ell]}^{(t)},\mathbf{P}_{[\ell]}^{(t)},\mathbf{Q}_{[\ell]}^{(t)},\mathbf{R}_{[\ell]}^{(t)},v_{[\ell]}^{(t)}$
Algorithm 3 Tensorized Federated Heat-Kernel Enhanced Multi-View Clustering (Local FedHK-Tucker)
Input: Local tensor data $\mathcal{X}_{[\ell]}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{U}_{g},\mathbf{V}_{g},\mathbf{W}_{g}\}$, personalization parameters $\{\lambda_{[\ell]},\rho_{[\ell]}\}$, $g$ is the global model index
Initialize:
$\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{[\ell]}\mathcal{G}_{g}+(1-\lambda_{[\ell]})\mathcal{G}_{[\ell]}$
$\mathbf{U}_{[\ell]}^{(0)},\mathbf{V}_{[\ell]}^{(0)},\mathbf{W}_{[\ell]}^{(0)}\leftarrow$ similarly personalized initialization
$v_{[\ell]}^{(0)}\leftarrow\frac{1}{s(\ell)}\mathbf{1}$
$t\leftarrow 0$
repeat
  $t\leftarrow t+1$
  // Compute tensorized heat-kernel coefficients
  Calculate $\delta_{[\ell]ij}^{h}$ using Eq. 61
  // Reconstruct cluster centers from decomposition
  $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$
  // Update memberships
  Update $U_{[\ell]}^{*^{(t)}}$ using Eq. 79
  // Update tensor factors using ALS
  Update $\mathcal{G}_{[\ell]}^{(t)}$ using Eq. 80
  Update $\mathbf{U}_{[\ell]}^{(t)}$ using Eq. 83
  Update $\mathbf{V}_{[\ell]}^{(t)}$ using Eq. 84
  Update $\mathbf{W}_{[\ell]}^{(t)}$ using Eq. 85
  // Update view weights
  Update $v_{[\ell]}^{(t)}$ using Eq. 81
  // Check convergence
  Compute $J_{[\ell]}^{tensor}$ using Eq. 75
until convergence or max iterations
return $U_{[\ell]}^{*^{(t)}},\mathcal{G}_{[\ell]}^{(t)},\mathbf{U}_{[\ell]}^{(t)},\mathbf{V}_{[\ell]}^{(t)},\mathbf{W}_{[\ell]}^{(t)},v_{[\ell]}^{(t)}$
6.5 Federated Aggregation with Personalization
Personalization in federated learning allows clients to maintain individualized models while benefiting from shared knowledge. In the context of heat-kernel enhanced multi-view clustering, personalization enables clients to adapt global clustering insights to their unique data distributions. To enable personalization, we introduce client-specific parameters that modulate the influence of global model updates on local clustering results.
6.5.1 Global Objective with Personalization
This section extends the federated aggregation framework to incorporate personalization for the heat-kernel enhanced multi-view clustering. First, we define the global objective function that balances local clustering quality with personalization constraints. The global model maintains shared cluster centers and view weights, while clients adapt these based on their local data characteristics.
The global model parameters are defined as:
  • Global cluster centers: $A_{\text{global}}=\{a_{kj}^{h,(\text{global})}\}_{k=1,j=1,h=1}^{c,d_{h},s}$
  • Global view weights: $V_{\text{global}}=\{v_{h}^{(\text{global})}\}_{h=1}^{s}$
The global federated objective with personalization is formulated as follows.
Definition 1 (Federated Heat-Kernel Enhanced Multi-View Clustering with Personalization).
The global federated objective balances local clustering quality with personalization constraints:
$J_{\text{global}}=\sum_{\ell=1}^{M}\left[\omega_{\ell}J_{FedHK}^{(\ell)}(U_{[\ell]},A_{[\ell]},V_{[\ell]})+\gamma\mathcal{R}_{person}(A_{[\ell]},A_{\text{global}})+\eta\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}}\right]$ (86)
where:
  • $\omega_{\ell}$ represents client-specific weights based on data quality and quantity
  • $\mathcal{R}_{person}(A_{[\ell]},A_{\text{global}}=\sum_{h,k}\|A_{[\ell]}^{h}-A_{\text{global}}^{h}\|_{F}^{2}$ ensures coordination between local and global centers
  • $\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})=\|V_{[\ell]}-V_{\text{global}}\|_{2}^{2}$ maintains view weight consistency
  • $\gamma,\eta$ control personalization vs. coordination trade-offs
We call this framework Federated Heat-Kernel Enhanced Multi-View Clustering with Personalization (FedHK-MVC-Person). Simply recall the local objective from FedHK-PARAFAC2 in Eq. 64 and FedHK-Tucker in Eq. 75 to instantiate $J_{FedHK}^{(\ell)}$ accordingly.
6.5.2 Privacy-Preserving Aggregation
To maintain privacy, clients share only aggregated statistics rather than raw cluster assignments:
Definition 2 (Federated Aggregation Protocol).
Each client $\ell$ computes and shares:
$S_{[\ell]}^{centers}$ $=\left\{\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}x_{[\ell]ij}^{h},\sum_{j=1}^{n_{i}}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\right\}_{kj,h}$ (87)
$S_{[\ell]}^{views}$ $=\left\{\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c}(\mu_{[\ell]ik}^{*})^{m}\text{KED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})\right\}_{h}$ (88)
$S_{[\ell]}^{quality}$ $=J_{[\ell]}\text{ and }n(\ell)$ (89)
The server performs secure aggregation:
$A_{\text{global}}^{h,(t+1)}=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{\ell}^{centers,h}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}\sum_{j=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}}$ (90)
$V_{\text{global}}^{(t+1)}=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)V_{\ell}^{(t)}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (91)
where client weights are computed as:
$\omega_{\ell}^{(t)}=\frac{\exp(-\tau J_{[\ell]}^{(t)})}{\sum_{\ell=1}^{M}\exp(-\tau J_{[\ell]}^{(t)})}$ (92)
6.5.3 Personalized Update Mechanism
Clients receive global updates and perform personalized integration:
$A_{\ell}^{h,(t+1)}$ $=\lambda_{\ell}A_{\ell}^{h,(t)}+(1-\lambda_{\ell})A_{\text{global}}^{h,(t+1)}$ (93)
$V_{\ell}^{(t+1)}$ $=\rho_{\ell}V_{\ell}^{(t)}+(1-\rho_{\ell})V_{\text{global}}^{(t+1)}$ (94)
where $\lambda_{\ell},\rho_{\ell}\in[0,1]$ are client-specific personalization parameters that can be learned adaptively:
$\lambda_{\ell}^{(t+1)}=\lambda_{\ell}^{(t)}-\beta\nabla_{\lambda_{\ell}}J_{[\ell]}(\lambda_{\ell}A_{\ell}^{(t)}+(1-\lambda_{\ell})A_{\text{global}}^{(t)})$ (95)
6.6 Tensorized Federated Aggregation with Personalization
The federated learning framework coordinates local tensorized heat-kernel enhanced clustering results while preserving client personalization. The aggregation process operates on tensor decomposition factors: core tensors, factor matrices, and view importance weights.
6.6.1 Tensorized Global Objective with Personalization
The tensorized global federated objective balances local clustering quality with personalization constraints:
$J_{\text{tensor(global)}}=\sum_{\ell=1}^{M}\left[\omega_{\ell}J_{[\ell]}^{tensor}(\mathcal{U}_{[\ell]},\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]},\mathbf{V}_{[\ell]})+\gamma\mathcal{R}_{tensor}(\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]})+\eta\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})\right]$ (96)
where the tensorized regularization terms are:
$\mathcal{R}_{tensor}$ $=\|\mathcal{G}_{[\ell]}-\mathcal{G}_{\text{global}}\|_{F}^{2}+\|\mathbf{P}_{[\ell]}-\mathbf{P}_{\text{global}}\|_{F}^{2}+\|\mathbf{Q}_{[\ell]}-\mathbf{Q}_{\text{global}}\|_{F}^{2}+\|\mathbf{R}_{[\ell]}-\mathbf{R}_{\text{global}}\|_{F}^{2}$ (97)
$\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})$ $=\|V_{[\ell]}-V_{\text{global}}\|_{2}^{2}$ (98)
6.6.2 Privacy-Preserving Tensor Aggregation
To maintain privacy while enabling effective tensor factor coordination, clients share only aggregated tensor statistics rather than raw factors:
Definition 3 (Tensorized Federated Aggregation Protocol).
Each client $\ell$ computes and shares:
$S_{[\ell]}^{core}$ $=\left\{\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c}(\mu_{[\ell]ik}^{*})^{m}\mathcal{G}_{[\ell]i,k,:,:}\right\}$ (99)
$S_{[\ell]}^{factors}$ $=\left\{\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\mathbf{P}_{[\ell]i,j,:},\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\mathbf{Q}_{[\ell]i,:,h},\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ij}^{*})^{m}\mathbf{R}_{[\ell]i,:,h}\right\}$ (100)
$S_{[\ell]}^{views}$ $=\left\{\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}_{[\ell]}(i,k,h)\right\}_{h}$ (101)
$S_{[\ell]}^{quality}$ $=J_{[\ell]}^{tensor}\text{ and }n(\ell)$ (102)
The server performs secure tensor aggregation:
$\mathcal{G}_{\text{global}}^{(t+1)}$ $=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{[\ell]}^{core}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (103)
$\mathbf{P}_{\text{global}}^{(t+1)}$ $=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{[\ell]}^{factors,P}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (104)
$\mathbf{Q}_{\text{global}}^{(t+1)}$ $=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{[\ell]}^{factors,Q}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (105)
$\mathbf{R}_{\text{global}}^{(t+1)}$ $=\frac{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}S_{[\ell]}^{factors,R}}{\sum_{\ell=1}^{M}\omega_{\ell}^{(t)}n(\ell)}$ (106)
6.6.3 Tensorized Personalized Update Mechanism
Clients receive global tensor updates and perform personalized integration across all tensor components:
$\mathcal{G}_{[\ell]}^{(t+1)}$ $=\lambda_{\ell}^{core}\mathcal{G}_{[\ell]}^{(t)}+(1-\lambda_{\ell}^{core})\mathcal{G}_{\text{global}}^{(t+1)}$ (107)
$\mathbf{P}_{[\ell]}^{(t+1)}$ $=\lambda_{\ell}^{P}\mathbf{P}_{[\ell]}^{(t)}+(1-\lambda_{\ell}^{P})\mathbf{P}_{\text{global}}^{(t+1)}$ (108)
$\mathbf{Q}_{[\ell]}^{(t+1)}$ $=\lambda_{\ell}^{Q}\mathbf{Q}_{[\ell]}^{(t)}+(1-\lambda_{\ell}^{Q})\mathbf{Q}_{\text{global}}^{(t+1)}$ (109)
$\mathbf{R}_{[\ell]}^{(t+1)}$ $=\lambda_{\ell}^{R}\mathbf{R}_{[\ell]}^{(t)}+(1-\lambda_{\ell}^{R})\mathbf{R}_{\text{global}}^{(t+1)}$ (110)
where $\lambda_{\ell}^{core},\lambda_{\ell}^{P},\lambda_{\ell}^{Q},\lambda_{\ell}^{R},\rho_{\ell}\in[0,1]$ are component-specific personalization parameters.
Algorithm 4 provides a comprehensive framework for implementing personalized FedHK-PARAFAC2. In order to achieve a personalized FedHK-Tucker, two changes must be made. Initially, adjustments are made to the tensor factor implication stage and its localization tensorized HKC. Secondly, modifications are made to the tensor aggregation stage through the implementation of Tucker decomposition.
Algorithm 4 Personalized FedHK-PARAFAC2
Server Initialization:
 Initialize global tensor factors $\mathcal{G}_{\text{global}}^{(0)},\mathbf{P}_{\text{global}}^{(0)},\mathbf{Q}_{\text{global}}^{(0)},\mathbf{R}_{\text{global}}^{(0)}$ and view weights $V_{\text{global}}^{(0)}$
 Set communication rounds $T$, local iterations $E$, tensor ranks $\{r_{2},r_{3}\}$
for $t=0,1,\ldots,T-1$ do
  Broadcast: Send $\mathcal{G}_{\text{global}}^{(t)},\mathbf{P}_{\text{global}}^{(t)},\mathbf{Q}_{\text{global}}^{(t)},\mathbf{R}_{\text{global}}^{(t)},V_{\text{global}}^{(t)}$ to all clients
  Local Computation (in parallel for all clients):
  for each client $\ell\in\{1,\ldots,M\}$ do
   // Tensorize local multi-view data
   Organize $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ into tensor $\mathcal{X}_{[\ell]}$
   // Personalized tensor factor initialization
   $\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{core}\mathcal{G}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{core})\mathcal{G}_{[\ell]}$
   $\mathbf{P}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{P}\mathbf{P}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{P})\mathbf{P}_{[\ell]}$
   $\mathbf{Q}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{Q}\mathbf{Q}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{Q})\mathbf{Q}_{\ell}$
   $\mathbf{R}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{R}\mathbf{R}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{R})\mathbf{R}_{[\ell]}$
   $A_{[\ell]}^{(0)}\leftarrow\rho_{\ell}A_{\text{global}}^{(t)}+(1-\rho_{\ell})A_{[\ell]}$
   $V_{[\ell]}^{(0)}\leftarrow\rho_{\ell}V_{\text{global}}^{(t)}+(1-\rho_{\ell})V_{[\ell]}$
   // Local tensorized heat-kernel enhanced clustering
   for $e=1,2,\ldots,E+1$ do
    Compute $\delta_{[\ell]ij}^{h}$ using Eq. 61
    Reconstruct $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$
    Perform Algorithm 2
   end for
   // Compute tensorized aggregation statistics
   Compute $S_{[\ell]}^{core},S_{[\ell]}^{factors},S_{[\ell]}^{views},S_{[\ell]}^{quality}$
   Send statistics to server
  end for
  Server Tensor Aggregation:
  Compute client weights $\omega_{\ell}$ using tensor quality measures
  Update $\mathcal{G}_{\text{global}}^{(t+1)}$ using Eq. 103
  Update $\mathbf{P}_{\text{global}}^{(t+1)}$ using Eq. 104
  Update $\mathbf{Q}_{\text{global}}^{(t+1)}$ using Eq. 105
  Update $\mathbf{R}_{\text{global}}^{(t+1)}$ using Eq. 106
  Update $A_{\text{global}}^{(t+1)}$ using Eq. 90
  Update $V_{\text{global}}^{(t+1)}$ using Eq. 91
  Adaptive Personalization:
  Update tensor-specific personalization parameters $\{\lambda_{\ell}^{core},\lambda_{\ell}^{P},\lambda_{\ell}^{Q},\lambda_{\ell}^{R},\rho_{\ell}\}$
end for
return Global tensor model $(\mathcal{G}_{\text{global}}^{(T)},\mathbf{P}_{\text{global}}^{(T)},\mathbf{Q}_{\text{global}}^{(T)},\mathbf{R}_{\text{global}}^{(T)},A_{\text{global}}^{(T)},V_{\text{global}}^{(T)})$ and personalized models $\{(\mathcal{G}_{[\ell]}^{(T)},\mathbf{P}_{[\ell]}^{(T)},\mathbf{Q}_{[\ell]}^{(T)},\mathbf{R}_{[\ell]}^{(T)},A_{[\ell]}^{(T)},V_{[\ell]}^{(T)})\}_{\ell=1}^{M}$
6.7 Theoretical Analysis
6.7.1 Convergence Properties
Theorem 4 (Local Convergence).
For each client $\ell$, the local heat-kernel enhanced multi-view fuzzy clustering algorithm converges to a local minimum of the objective function $J_{[\ell]}$.
Proof.
The proof follows from the monotonic decrease property of alternating optimization. Each update step:
  1. 1.
    Membership updates (Eq. 26) minimize $J_{[\ell]}$ w.r.t. $U_{[\ell]}^{*}$ for fixed $A_{[\ell]},V_{[\ell]}$
  2. 2.
    Center updates (Eq. 27) minimize $J_{[\ell]}$ w.r.t. $A_{[\ell]}$ for fixed $U_{[\ell]},V_{[\ell]}$
  3. 3.
    View weight updates (Eq. 28) minimize $J_{[\ell]}$ w.r.t. $V_{[\ell]}$ for fixed $U_{[\ell]},A_{[\ell]}$
Since $J_{[\ell]}$ is bounded below by zero and decreases monotonically, convergence is guaranteed. ∎
Theorem 5 (Global Convergence).
Under mild regularity conditions, the federated algorithm converges to a stationary point of the global objective function $J_{\text{global}}$.
Proof.
The global convergence follows from the convergence theory of federated optimization. The key insight is that each local update corresponds to a proximal gradient step on the global objective with personalization regularization, ensuring convergence to a stationary point. ∎
6.7.2 Privacy Guarantees
Theorem 6 (Differential Privacy).
The federated aggregation protocol provides $(\epsilon,\delta)$-differential privacy when clients add calibrated noise to their shared statistics.
Proof.
The shared statistics $S_{[\ell]}^{centers},S_{[\ell]}^{views},S_{[\ell]}^{quality}$ have bounded sensitivity. Adding Gaussian noise with variance $\sigma^{2}=\frac{2\Delta^{2}\log(1.25/\delta)}{\epsilon^{2}}$ where $\Delta$ is the sensitivity, ensures $(\epsilon,\delta)$-differential privacy. ∎
Proposition 1 (Communication Efficiency).
The federated protocol requires $\mathcal{O}(c(\ell)\cdot\bar{d}_{h}^{(\ell)}\cdot\bar{s}(\ell))$ communication per round, where $\bar{d}_{h}^{(\ell)}$ and $\bar{s}(\ell)$ are average feature dimensions and views across clients.
6.7.3 Computational Complexity
Theorem 7 (Time Complexity).
The local clustering complexity per client $\ell$ per iteration is $\mathcal{O}(n(\ell)c(\ell)s(\ell)D(\ell))$ where $D(\ell)=\sum_{h=1}^{s(\ell)}d_{h}^{(\ell)}$ is the total local dimensionality.
The federated approach maintains the same asymptotic complexity as centralized multi-view clustering while providing personalization and privacy benefits.
6.7.4 Tensorized Complexity Analysis
Theorem 8 (Tensorized Time Complexity).
The tensorized local clustering complexity per client $\ell$ per iteration is $\mathcal{O}(n(\ell)c(\ell)r_{2}r_{3}+n(\ell)cD(\ell)s(\ell))$ where $r_{2},r_{3}$ are the tensor decomposition ranks.
Proof.
The complexity breakdown includes:
  1. 1.
    Heat-kernel coefficient computation: $\mathcal{O}(n(\ell)D(\ell)s(\ell))$
  2. 2.
    Tensor reconstruction: $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)D(\ell)r_{2}+c(\ell)s(\ell)r_{3})$
  3. 3.
    Membership updates: $\mathcal{O}(n(\ell)c(\ell)s(\ell)D(\ell))$
  4. 4.
    Tucker decomposition updates: $\mathcal{O}(n(\ell)c(\ell)r_{2}r_{3})$ per factor using ALS
The dominant term is $\mathcal{O}(n(\ell)c(\ell)D(\ell)s(\ell))$ when $r_{2},r_{3}\ll\min(D(\ell),s(\ell))$. ∎
Theorem 9 (Tensorized Communication Efficiency).
The tensorized federated protocol requires $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$ communication per round, compared to $\mathcal{O}(c(\ell)\bar{D}(\ell)\bar{s}(\ell))$ for the non-tensorized version.
The tensorized approach provides significant communication savings when $r_{2}\ll\bar{D}(\ell)$ and $r_{3}\ll\bar{s}(\ell)$, which is typically the case in practice.
Lemma 1 (Tensorized Approximation Quality).
The reconstruction error of the tensorized cluster centers is bounded by:
$\|\mathcal{A}_{[\ell]}-\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}\|_{F}^{2}\leq\epsilon_{tensor}$ (111)
where $\epsilon_{tensor}$ depends on the chosen tensor ranks and the intrinsic tensor rank of the true cluster center tensor.
6.7.5 Personalization-Accuracy Trade-off
Lemma 2 (Personalization Bound).
The personalization error is bounded by:
$\|A_{[\ell]}-A_{\text{global}}\|_{F}^{2}\leq\frac{\lambda_{\ell}^{2}}{1-\lambda_{\ell}^{2}}\|A_{[\ell]}-A_{\text{global}}\|_{F}^{2}$ (112)
where $A_{[\ell]}$ is the purely local optimum.
This bound shows that personalization parameters $\lambda_{\ell}$ control the trade-off between local specialization and global coordination.
7 Conclusion
The present paper proposes a novel personalized federated learning framework that integrates heat-kernel enhanced multi-view clustering with tensor-based global coordination. The primary contributions of this study include the introduction of a novel method, Heat-kernel Enhanced Federated Clustering, which has not been previously documented. The adoption of heat-kernel coefficients in federated multi-view clustering facilitates the more effective capture of complex local patterns while ensuring the maintenance of global coordination. In the Personalized Aggregation Mechanism, the framework provides an adaptive personalization that balances local specialization with global knowledge sharing through learnable personalization parameters. In the context of Privacy-Preserving Protocol, the aggregation protocol has been engineered to ensure differential privacy while preserving the efficacy of clustering through the implementation of meticulously designed statistical sharing methodologies. A comprehensive theoretical analysis is provided, offering convergence guarantees, privacy bounds, and complexity analysis. These elements serve to establish the theoretical foundation of the approach.
7.1 Future Directions
This work suggests several promising research directions. Initially, the Dynamic View Discovery method was employed. The framework is being extended to accommodate the dynamic alteration of view structures during federated rounds. The present study explores the potential of hierarchical federated clustering for large-scale deployment, investigating multi-level federated architectures in the process. In the context of adversarial robustness, the focus is on the analysis and enhancement of systems’ resilience to malicious clients and data poisoning attacks. The concept of continuous learning involves the adaptation of a framework for federated continual learning scenarios, with the understanding that data distributions undergo evolution over time. In the domain of cross-modal applications, the focus is on the exploration of applications to federated learning with fundamentally different data modalities, including but not limited to text, images, and sensors. The proposed framework presents novel opportunities for privacy-preserving, personalized machine learning on distributed multi-view data, with extensive applications in healthcare, the Internet of Things (IoT), and collaborative intelligence systems. The proposed framework opens new possibilities for privacy-preserving, personalized machine learning on distributed multi-view data, with broad applications in healthcare, IoT, and collaborative intelligence systems.
7.2 Potential Limitations
The proposed frameworks, while innovative, present several potential limitations. The computational complexity associated with tensor decompositions may pose challenges for clients with limited resources. The selection of personalization parameters requires careful tuning to balance local adaptation and global coordination effectively. The assumption of data homogeneity across clients may not hold in all scenarios, potentially impacting clustering performance. Additionally, the privacy guarantees depend on the proper calibration of noise, which may affect the utility of the shared statistics. Future work will focus on addressing these limitations through algorithmic optimizations, adaptive parameter selection methods, and robust privacy-preserving techniques with minimal utility loss and enhanced adaptability to heterogeneous data distributions. Validations both on synthetic and real-world datasets will be conducted to assess the practical performance and limitations of the proposed frameworks.
7.3 Applications and Implications
The proposed personalized federated heat-kernel enhanced multi-view clustering framework has significant applications across various domains. In healthcare, it can facilitate collaborative analysis of patient data from multiple institutions while preserving privacy. In IoT, the framework can enable efficient clustering of sensor data collected from distributed devices, enhancing real-time decision-making. In collaborative intelligence systems, it can support the integration of diverse data sources for improved insights and predictions. The implications of this work extend to advancing privacy-preserving machine learning techniques, promoting personalized models that respect individual client characteristics, and fostering collaboration in distributed environments without compromising data security. With the aggregation of several factors such as privacy, personalization, and multi-view data integration using tensorization, this framework sets a new standard for federated learning applications in complex, real-world scenarios.
Acknowledgment: This paper was authored during the period of the author’s postdoctoral studies at ISTI-CNR, Italy.
Funding: This research received no external funding.
Conflicts of interest: The author declare no conflict of interest.